### v1

# An Elementary Approach to Scheduling in Generative Diffusion Models

Abstract
An elementary approach to characterizing the impact of noise scheduling and time discretization in generative diffusion models is developed. Considering a simplified model where the source distribution is multivariate Gaussian with a given covariance matrix, the explicit closed-form evolution trajectory of the distributions across reverse sampling steps is derived, and consequently, the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output is obtained. The effect of the number of time discretization steps on the convergence of this KL divergence is studied via the Euler-Maclaurin expansion. An optimization problem is formulated, and its solution noise schedule is obtained via calculus of variations, shown to follow a tangent law whose coefficient is determined by the eigenvalues of the source covariance matrix. For an alternative scenario, more realistic in practice, where pretrained models have been obtained for some given noise schedules, the KL divergence also provides a measure to compare different time discretization strategies in reverse sampling. Experiments across different datasets and pretrained models demonstrate that the time discretization strategy selected by our approach consistently outperforms baseline and search-based strategies, particularly when the budget on the number of function evaluations is very tight.
I Introduction
Diffusion models (DMs) [18, 5, 20] have established themselves as a dominant force in generative modeling, delivering state-of-the-art results across diverse applications. The fundamental mechanism of DMs involves the learning of the conditional distribution of a reverse sampling process through a predefined noise-adding forward process.
Research on the forward diffusion process typically focuses on optimizing the noise schedule, prediction target, and network architecture to yield a superior estimator. While recent studies have identified effective configurations via empirical testing or Fisher information analysis [8, 6, 3, 17, 26], these analyses generally fall short of directly assessing the impact of the noise schedule on the accuracy of the distribution generated by the reverse sampling process. Regarding this reverse sampling process, although the paradigm has shifted from stochastic differential equation (SDE) to more efficient deterministic ordinary differential equation (ODE) solvers [19, 12, 13, 27, 28], the generation procedure still necessitates multiple forward evaluations of a large network, with the cost measured by the number of function evaluations (NFEs), to substantiate its exceptional capabilities.
To improve generation quality under a fixed budget of NFEs, optimizing the time discretization strategy is crucial [19, 12, 6]. Recognizing this, a growing body of research has been conducted [22, 10, 16, 25, 23, 24, 29]. However, the majority of extant works rely on computationally expensive data-driven retraining or evaluation processes for searching the optimal steps, typically incurring a substantial overhead.
Aiming at understanding from first principles, the interplay between noise scheduling and sampling efficiency, in this work, we adopt an elementary approach based on an idealized case where the source distribution is multivariate Gaussian with a given covariance matrix. Such a simple model induces a linear form of the optimal posterior estimator, and consequently an explicit closed-form characterization of the evolution trajectory of the distributions across reverse sampling steps. Using this closed-form formulation, we can conveniently investigate the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output.
We demonstrate two applications of this KL divergence-based analytical tool. In the first application, we study how fast the KL divergence asymptotically vanishes as the number of time discretization steps increases. Adopting the Euler-Maclaurin expansion [30], we obtain the precise asymptotic behavior of the reverse sampling process and identify the component that dominates the KL divergence. These allow us to establish a direct connection between the noise schedule and the sampling accuracy. Based on this, we propose and solve a variational optimization problem that yields a tangent law of the noise schedule, whose coefficient depends upon the eigenvalues of the source covariance matrix.
In the second application, we consider the practically relevant scenario where a DM has already been trained, under some prescribed noise schedule. We can then employ the KL divergence as a measure to compare different time discretization strategies in the reverse sampling process. This provides a principled low-cost approach to selecting efficient time discretization designs, without requiring any additional retraining or search. We test this approach over CIFAR-10 [9] and FFHQ-64 [7] datasets with different model configurations as outlined in [6, 20], and the experimental results confirm that our strategy consistently demonstrates superior performance in comparison to baseline and state-of-the-art search-based strategies, particularly when the budget on NFEs is very tight.
Notation: For brevity, we abbreviate a function $f(t)$ as $f$ and denote its discrete-time counterpart $f(t_{j})$ by $f_{t_{j}}$. Derivatives with respect to $t$ are written as $\dot{f}=\mathrm{d}f/\mathrm{d}t$ and $\ddot{f}=\mathrm{d}^{2}f/\mathrm{d}t^{2}$. Throughout $\log(\cdot)$ represents the natural logarithm.
II Preliminaries
In this section, we first review the mechanism of DMs, encompassing the forward diffusion process with its noise schedule design, and the deterministic reverse sampling process employed for generation. Subsequently, we outline the multivariate Gaussian setting that enables a tractable theoretical analysis.
II-A Diffusion Models
The forward diffusion process of DMs is characterized by a linear Gaussian perturbation [5, 8]:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
where $t\in[0,1]$ denotes the diffusion time, $\mathbf{x}_{0}\sim q(\mathbf{x})$ represents the source data sample, and $\alpha_{t},\sigma_{t}$ parameterize the noise schedule that directly controls the noise-adding process.
Commonly adopted noise schedules include two types: variance-preserving (VP) and variance-exploding (VE) [20]. The VP setting usually satisfies the constraint $\alpha^{2}(t)+\sigma^{2}(t)=1$, where $\alpha(t)$ decays from 1 to 0. In contrast, the VE setting maintains $\alpha(t)\equiv 1$, while $\sigma(t)$ increases from 0 to a predefined maximum $\sigma_{\max}\gg 1$. In either case, the forward transition kernel is given by
$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}\!\left(\alpha_{t}\mathbf{x}_{0},\,\sigma^{2}_{t}\mathbf{I}\right),$ (2)
which progressively diffuses the source distribution $q(\mathbf{x}_{0})$ towards a Gaussian noise as $t\to 1$.
The deterministic reverse sampling process, which reverses the forward diffusion process (1) to generate data samples from $q(\mathbf{x})$, is governed by the following update formula [19, 12], for each step $j\in\left[1,N\right]$:
$\hat{\mathbf{x}}_{t_{j-1}}=\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\hat{\mathbf{x}}_{t_{j}}+\left(\sigma_{t_{j-1}}-\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\sigma_{t_{j}}\right)\bm{\epsilon}_{\theta}(\hat{\mathbf{x}}_{t_{j}},t_{j}),$ (3)
where $\{t_{i}\}_{i=0}^{N}$ denotes the time discretization sequence decreasing from $t_{N}=1$ to $t_{0}=0$, and $\bm{\epsilon}_{\theta}$ is a learned noise estimator. The reverse sampling process is initialized with $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ for the VP setting and $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{\max}^{2}\mathbf{I})$ for the VE setting.
To obtain the noise estimator $\bm{\epsilon}_{\theta}$, the training objective is to minimize the following mean squared error [5]:
$\mathcal{L}_{\theta}=\mathbb{E}_{t\in\left[0,1\right],\mathbf{x}_{0}\sim q(\mathbf{x}),\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)}\!\left[\bigl\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\bigr\|^{2}\right],$ (4)
which implies that the optimal estimator is the conditional expectation, i.e., $\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=\mathbb{E}[\bm{\epsilon}|\mathbf{x}_{t}]$. Inverting the linear forward relation in (1), then gives the identity:
$\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=\mathbb{E}\left[\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}}\bigg|\mathbf{x}_{t}\right]=\frac{\mathbf{x}_{t}-\alpha_{t}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\sigma_{t}}.$ (5)
Thus, learning to predict the noise $\bm{\epsilon}$ is mathematically equivalent to estimating the posterior mean of the source data $\mathbf{x}_{0}$.
II-B Gaussian Setup
To facilitate theoretical analysis, we consider the idealized case where the source distribution is multivariate Gaussian with a given covariance matrix, i.e., $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{\mathbf{x}})$. Without loss of generality, we assume that the covariance matrix $\bm{\Sigma}_{\mathbf{x}}$ is positive definite. Let the eigendecomposition of the covariance matrix be $\bm{\Sigma}_{\mathbf{x}}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{-1}$, where $\mathbf{U}$ is an orthogonal matrix and $\mathbf{\Lambda}=\text{diag}(\mu_{1},\ldots,\mu_{k})$ denotes the diagonal matrix of eigenvalues. Furthermore, due to the linearity of the forward diffusion process (1), the marginal distribution $q(\mathbf{x}_{t})$ remains Gaussian:
$q(\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{0},\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right).$ (6)
Moreover, the optimal posterior estimator admits a closed-form solution [15]:
$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\alpha_{t}\bm{\Sigma}_{\mathbf{x}}\left(\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right)^{-1}\mathbf{x}_{t}.$ (7)
III KL-Based Analysis and Applications
In this section, we present the main results derived from our theoretical analysis under the Gaussian setting, and demonstrate their applications. For clarity of exposition, we focus on the VP setting; analogous results for the VE setting are briefly described in discussions and remarks.
III-A KL Divergence Analysis
We begin by characterizing the evolution trajectory of the distributions across reverse sampling steps, as given by the following lemma.
Lemma 1.
With the initialization $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, time discretization sequence $\{t_{i}\}_{i=0}^{N}$, and the optimal estimator given by (5) and (7), the generated sample $\hat{\mathbf{x}}_{t_{0}}$ produced by the reverse sampling process (3) is distributed as
$p\left(\hat{\mathbf{x}}_{t_{0}}\right)=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),\quad\mathbf{M}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (8)
where the $\ell$-th eigenvalue is given by
$m_{\ell}=\prod_{j=1}^{N}\left(\frac{\alpha_{t_{j-1}}\alpha_{t_{j}}\mu_{\ell}+\sigma_{t_{j-1}}\sigma_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\right)^{2},\quad\ell\in\{1,\ldots,k\}.$ (9)
Proof.
See Appendix A. ∎
By (6), we can rewrite $q(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$, where $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$ with
$n_{\ell}\triangleq\alpha_{t_{0}}^{2}\mu_{\ell}+\sigma_{t_{0}}^{2}.$ (10)
According to Lemma 1, the KL divergence is expressible in closed form:
$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)=\frac{1}{2}\sum_{\ell=1}^{k}\left(\frac{m_{\ell}}{n_{\ell}}-\log\frac{m_{\ell}}{n_{\ell}}-1\right),$ (11)
by the standard KL divergence expression between two Gaussian distributions.
For the sake of the subsequent analysis, for each $\mu_{\ell}$, we define
$S_{\ell}^{N}\triangleq\sum_{j=1}^{N}\log\!\left(\frac{\alpha_{t_{j-1}}\alpha_{t_{j}}\mu_{\ell}+\sigma_{t_{j-1}}\sigma_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\right),$ (12)
so that, from (9) we have $S_{\ell}^{N}=\frac{1}{2}\log m_{\ell}.$
To further investigate the asymptotic convergence behavior of the KL divergence (11) with respect to the number of time discretization steps $N$, we present the following lemma derived by the Euler-Maclaurin expansion.
Lemma 2.
Consider $S_{\ell}^{N}$ defined in (12) with the uniform discretization $t_{j}\triangleq j/N$ on $[0,1]$. Assume the schedule functions $\alpha,\sigma$ are sufficiently smooth. Define
$I_{\ell}\triangleq\int_{0}^{1}F_{\ell}(t)\,\mathrm{d}t,\quad F_{\ell}(t)\triangleq-\frac{\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma}}{\alpha^{2}\mu_{\ell}+\sigma^{2}}.$ (13)
Then, as $N\to\infty$, we obtain
$S_{\ell}^{N}=I_{\ell}+\frac{E_{\ell}^{1}}{N}+\mathcal{O}\!\left(\frac{1}{N^{2}}\right),$ (14)
where $E_{\ell}^{1}$ is independent of $N$, specifically,
$E_{\ell}^{1}=-\frac{\mu_{\ell}}{2}\int_{0}^{1}\frac{(\alpha\dot{\sigma}-\sigma\dot{\alpha})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t.$ (15)
Moreover, it holds that
$I_{\ell}=\frac{1}{2}\log n_{\ell}.$ (16)
Proof.
See Appendix B. ∎
Combining the definition (12) with the closed-form solution (16) and the expansion (14) in Lemma 2, we can explicitly express the residual $r_{\ell}$ as
$r_{\ell}\triangleq S_{\ell}^{N}-I_{\ell}=\frac{1}{2}\log\frac{m_{\ell}}{n_{\ell}}=\frac{E_{\ell}^{1}}{N}+\mathcal{O}\!\left(\frac{1}{N^{2}}\right).$ (17)
This implies that $r_{\ell}=\mathcal{O}(1/N)$, which justifies the use of the Taylor expansion for large $N$. Substituting (17) into (11), the KL divergence can be expanded as follows:
$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)$ $=\frac{1}{2}\sum_{\ell=1}^{k}\bigl(e^{2r_{\ell}}-2r_{\ell}-1\bigr)$
$=\sum_{\ell=1}^{k}r_{\ell}^{2}+\mathcal{O}\!\left(\sum_{\ell=1}^{k}|r_{\ell}|^{3}\right)$
$=\frac{1}{N^{2}}\sum_{\ell=1}^{k}(E_{\ell}^{1})^{2}+\mathcal{O}\!\left(\frac{1}{N^{3}}\right).$ (18)
Here, the second equality follows from the Taylor expansion $e^{x}-x-1=\frac{1}{2}x^{2}+\mathcal{O}(|x|^{3})$ with $x=2r_{\ell}$, and the last equality results from substituting the leading order term of $r_{\ell}$.
Since the higher-order terms vanish faster as $N\to\infty$, minimizing the KL divergence at a fixed $N$ asymptotically boils down to minimizing the coefficient of the leading $1/N^{2}$ term. This yields the following objective functional subject to corresponding boundary conditions:
$\min_{\alpha,\,\sigma}\ \mathcal{L}[\alpha,\sigma]\triangleq\sum_{\ell=1}^{k}(E_{\ell}^{1})^{2}.$ (19)
Remark (Convergence and Consistency).
From (18), we observe that the KL divergence scales as $\mathcal{O}(1/N^{2})$. As $N\to\infty$, the eigenvalue mismatch vanishes (i.e., $m_{\ell}/n_{\ell}\to 1$), leading to $D_{\mathrm{KL}}(p(\hat{\mathbf{x}}_{t_{0}})\|q(\mathbf{x}_{t_{0}}))\to 0$. This verifies that the reverse sampling process in (3) is asymptotically consistent, recovering the exact source distribution in the continuous-time limit, in accordance with the classical theory well known for DMs (see, e.g., [1]).
III-B Derivation of Tangent Law
The preceding analysis of the KL divergence reveals a direct connection between the noise schedule and sampling accuracy, culminating in the formulation of the optimization problem (19). However, directly minimizing the objective functional in (19) over the schedule functions $\alpha(t)$ and $\sigma(t)$ is generally intractable. To gain analytical tractability, we decompose the problem and first analyze the optimality condition for a fixed $\mu_{\ell}$. This leads us to introduce a specific law of noise scheduling, defined as follows.
Definition 1.
For a fixed mode $\mu_{\ell}>0$, we define the tangent law schedule via the ratio $\eta(t)\triangleq\sigma(t)/\alpha(t)$ as
$\eta_{\ell}(t)=\sqrt{\mu_{\ell}}\tan\!\left(\frac{\pi}{2}t\right),\quad t\in[0,1).$ (20)
The following theorem establishes that this tangent law schedule is, in fact, the unique minimizer of the reverse sampling mismatch for the $\ell$-th mode.
Theorem 1.
The minimization of the term $(E_{\ell}^{1})^{2}$ in (19), with the expression of $E_{\ell}^{1}$ given in (15), is equivalent to minimizing
$\min_{\eta}\mathcal{J}_{\ell}[\eta]\triangleq\int_{0}^{1}\frac{\dot{\eta}^{2}}{\bigl(\mu_{\ell}+\eta^{2}\bigr)^{2}}\,\mathrm{d}t,$
$\text{subject to }\eta(0)=0\text{ and }\lim_{t\to 1}\eta(t)=\infty.$ (21)
The unique solution $\eta(\cdot)$ that minimizes the functional $\mathcal{J}_{\ell}$ is given by the tangent law schedule $\eta_{\ell}(t)$ in Definition 1.
Proof.
Using the quotient rule $\dot{\eta}=(\alpha\dot{\sigma}-\sigma\dot{\alpha})/\alpha^{2}$ and the relation $\alpha^{2}\mu_{\ell}+\sigma^{2}=\alpha^{2}(\mu_{\ell}+\eta^{2})$, (15) can be simplified to
$\frac{(\alpha\dot{\sigma}-\sigma\dot{\alpha})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}=\frac{\dot{\eta}^{2}}{(\mu_{\ell}+\eta^{2})^{2}},$
and with the VP setting, we have the boundary conditions (21). Thus, minimizing $(E_{\ell}^{1})^{2}$ is equivalent to minimizing $\mathcal{J}_{\ell}[\eta]$.
To solve this variational problem, we introduce a change of variables:
$Q_{\ell}(t)\triangleq\frac{1}{\sqrt{\mu_{\ell}}}\arctan\!\left(\frac{\eta(t)}{\sqrt{\mu_{\ell}}}\right).$ (22)
Differentiating $Q_{\ell}(t)$ with respect to $t$ yields
$\dot{Q}_{\ell}(t)=\frac{1}{\sqrt{\mu_{\ell}}}\cdot\frac{1}{1+(\eta/\sqrt{\mu_{\ell}})^{2}}\cdot\frac{\dot{\eta}}{\sqrt{\mu_{\ell}}}=\frac{\dot{\eta}}{\mu_{\ell}+\eta^{2}}.$ (23)
Consequently, the objective functional simplifies to $\mathcal{J}_{\ell}[\eta]=\int_{0}^{1}\bigl(\dot{Q}_{\ell}\bigr)^{2}\,\mathrm{d}t.$ The Euler–Lagrange equation [30] for the Lagrangian function $L(t,Q_{\ell},\dot{Q}_{\ell})=\dot{Q}_{\ell}^{2}$ is
$\frac{\partial L}{\partial Q_{\ell}}-\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial L}{\partial\dot{Q}_{\ell}}\right)=-2\ddot{Q}_{\ell}=0,$ (24)
implying that the optimal $Q_{\ell}(t)$ is linear in $t$: $Q_{\ell}(t)=ct+d$.
Given the boundary conditions $\eta(0)=0$ and $\lim_{t\to 1}\eta(t)=\infty$, the transformed boundary values are $Q_{\ell}(0)=0$ and $Q_{\ell}(1)=\frac{\pi}{2\sqrt{\mu_{\ell}}}$. Therefore, the optimal trajectory is $Q_{\ell}(t)=\frac{\pi t}{2\sqrt{\mu_{\ell}}}$. Finally, combining the definition of $Q_{\ell}(t)$ (22) yields the tangent law schedule $\eta_{\ell}(t)=\sqrt{\mu_{\ell}}\tan(\frac{\pi}{2}t)$. ∎
Theorem 1 establishes that the optimal schedule $\eta_{\ell}(t)$ depends explicitly on the specific eigenvalue $\mu_{\ell}$. However, practical diffusion models typically enforce a unified schedule shared across all data dimensions. To reconcile this theoretical optimality with practical scenarios, we introduce a global schedule parameterized by a scalar $\gamma>0$, adopting the form of tangent law derived in Theorem 1:
$\eta_{\gamma}(t)\triangleq\sqrt{\gamma}\,\tan\!\bigl(\frac{\pi}{2}t\bigr).$ (25)
Substituting this parameterized tangent law schedule into the leading order coefficient (15) and the global objective functional (19) yields the following choice of $\gamma$.
Theorem 2.
Under the parameterized tangent law schedule (25), the global objective functional (19) is strictly convex with respect to $\gamma$ and admits a unique closed-form minimizer,
$\gamma^{\star}=\sqrt{\frac{\sum_{\ell=1}^{k}\mu_{\ell}}{\sum_{\ell=1}^{k}\mu_{\ell}^{-1}}}=\sqrt{\frac{\mathrm{tr}(\bm{\Sigma}_{\mathbf{x}})}{\mathrm{tr}(\bm{\Sigma}_{\mathbf{x}}^{-1})}}.$ (26)
Proof.
As derived in Appendix D-A, the leading order coefficient (15) for the parameterized tangent law schedule (25) is given by
$E_{\ell}^{1}(\gamma)=-\frac{\pi^{2}}{16}\left(\sqrt{\frac{\mu_{\ell}}{\gamma}}+\sqrt{\frac{\gamma}{\mu_{\ell}}}\right).$ (27)
Substituting (27) into $\mathcal{L}(\gamma)=\sum_{\ell=1}^{k}\left(E_{\ell}^{1}\left(\gamma\right)\right)^{2}$ and omitting scaling factors and constant terms, we have that minimizing $\mathcal{L}(\gamma)$ is equivalent to minimizing the function
$J(\gamma)=\gamma\sum_{\ell=1}^{k}\mu_{\ell}^{-1}+\gamma^{-1}\sum_{\ell=1}^{k}\mu_{\ell}.$ (28)
The first and second derivatives of $J(\gamma)$ are
$J^{\prime}(\gamma)=\sum_{\ell=1}^{k}\mu_{\ell}^{-1}-\gamma^{-2}\sum_{\ell=1}^{k}\mu_{\ell},J^{\prime\prime}(\gamma)=2\gamma^{-3}\sum_{\ell=1}^{k}\mu_{\ell}.$ (29)
Since $\mu_{\ell}>0$, we have $J^{\prime\prime}(\gamma)>0$ for all $\gamma>0$, which implies $J(\gamma)$ is strictly convex. Finally, setting $J^{\prime}(\gamma)=0$ yields the unique minimizer in (26). ∎
(a) VP setting.
(b) VE setting.
Figure 1: KL divergence versus steps NN for the time discretization strategy (30) under (a) VP and (b) VE settings.
Remark (Optimization for the VE setting).
The derivation for the VE setting parallels the VP setting, but differs in the boundary condition at $t=1$. This introduces a dependency on the terminal noise level $\sigma_{\max}$ into the objective, which generally precludes a simple closed-form solution for $\gamma^{\star}$. Nevertheless, in the practical regime of $\sigma_{\max}^{2}\gg\mu_{\ell}$, the boundary effect becomes negligible. Thus, the VP optimizer $\gamma^{\star}$ in  (26) can still provide an accurate approximation (details in Appendix D-A).
Additional numerical results demonstrating that the parameterized tangent law schedule outperforms other noise schedules for the Gaussian setting are presented in Appendix D-B.
III-C Time Discretization Strategy Evaluation
In this subsection, we demonstrate another more direct application of the KL divergence-based analysis. Given that training a new DM is prohibitively computationally expensive for large datasets, in practice, users are typically constrained to pretrained models with prescribed noise schedules. Consequently, our focus shifts to designing time discretization strategies, as they crucially influence the generation quality of existing pretrained models. Specifically, we leverage the closed-form KL divergence (11) as a measure to assess different time discretization strategies. This enables us to identify effective time discretization designs without incurring significant computational overhead.
To be specific, let us first parameterize the steps in terms of the half-logSNR $\lambda(t)\triangleq\log(\alpha(t)/\sigma(t))$ and consider a power-uniform discretization companded with $\rho$ (using sign-preserving powers $sign(x)|x|^{\rho}$ when $x$ is negative):
$\lambda_{i}^{(\rho)}=\Big(\lambda_{t_{N}}^{1/\rho}+\tfrac{i}{N}\big(\lambda_{t_{0}}^{1/\rho}-\lambda_{t_{N}}^{1/\rho}\big)\Big)^{\rho},\quad i=0,\dots,N,$ (30)
which degenerates to the uniform-$\lambda$ rule introduced in [12] when $\rho=1$. With this, the time discretization steps used in the reverse sampling process (3) are recovered via $t_{i}=\lambda^{-1}(\lambda_{i}^{(\rho)})$, which only depends on the prescribed $\alpha(t),\sigma(t)$.
TABLE I: Quantitative comparison of time discretization strategies in terms of FID scores ($\downarrow$). Bold highlights the best result, while underlining marks the second best.
Sampler Step strategy NFEs
3 4 5 6 7 8 9 10 12 15 20 25 30 50 100
Model: EDM — Dataset: CIFAR-10
DPM-Solver++ Uniform-$\lambda$ 120.54 72.33 53.15 37.78 29.43 23.78 19.29 16.29 12.16 8.74 6.01 4.69 3.96 2.86 2.32
Xue’24 89.59 56.29 37.94 26.07 21.65 18.26 18.99 15.87 11.89 8.67 6.00 4.70 3.98 2.86 2.33
Ours ($\rho=1.5$) 82.52 54.95 35.93 26.76 20.01 16.31 13.26 11.47 8.70 6.48 4.67 3.83 3.34 2.60 2.24
Ours ($\rho=2.0$) 86.27 53.82 37.26 27.04 20.74 16.75 13.75 11.78 9.04 6.66 4.81 3.92 3.42 2.63 2.25
UniPC Uniform-$\lambda$ 119.07 69.04 48.41 32.47 23.77 18.14 13.83 11.13 7.65 5.07 3.38 2.76 2.48 2.16 2.08
Xue’24 83.91 50.00 29.89 18.70 14.49 11.63 13.23 10.48 7.44 5.02 3.38 2.76 2.49 2.16 2.08
Ours ($\rho=1.5$) 79.62 49.44 29.50 20.16 14.11 10.78 8.37 6.94 5.09 3.76 2.84 2.49 2.32 2.13 2.08
Ours ($\rho=2.0$) 78.88 45.79 28.27 18.94 13.84 10.72 8.67 7.22 5.45 4.08 3.06 2.63 2.42 2.16 2.08
Model: EDM — Dataset: FFHQ-64
DPM-Solver++ Uniform-$\lambda$ 88.98 63.98 50.69 40.48 32.92 27.63 23.69 20.61 16.24 12.28 8.79 6.97 5.90 4.10 3.11
Xue’24 89.62 55.70 41.09 29.80 25.77 22.13 22.55 20.33 15.99 12.14 8.75 6.98 5.89 4.10 3.12
Ours ($\rho=1.5$) 78.71 50.79 37.54 30.46 24.68 20.95 17.64 15.58 12.34 9.46 6.95 5.64 4.89 3.61 2.91
Ours ($\rho=2.0$) 82.31 54.54 38.60 29.96 24.75 21.18 18.07 15.89 12.63 9.67 7.09 5.75 4.95 3.63 2.90
UniPC Uniform-$\lambda$ 88.15 61.85 47.19 35.91 27.71 22.04 17.93 14.85 10.77 7.47 5.02 3.98 3.46 2.78 2.53
Xue’24 83.65 49.87 33.37 22.35 18.38 14.99 16.26 14.25 10.51 7.36 4.99 3.99 3.47 2.79 2.53
Ours ($\rho=1.5$) 75.76 46.25 31.69 23.76 18.14 14.41 11.62 9.80 7.32 5.35 3.91 3.29 2.99 2.61 2.49
Ours ($\rho=2.0$) 75.49 46.84 30.18 21.66 17.13 14.05 11.73 10.01 7.68 5.69 4.12 3.42 3.05 2.61 2.48
Model: VP-SDE — Dataset: CIFAR-10
DPM-Solver++ Uniform-$\lambda$ 93.66 71.75 49.70 38.16 30.38 24.91 20.93 17.91 13.95 10.39 7.54 6.07 5.29 3.96 3.24
Xue’24 52.80 38.45 24.85 19.13 16.61 18.02 20.47 15.11 12.70 10.28 7.51 6.05 5.28 3.95 3.26
Ours ($\rho=1.5$) 53.56 41.23 33.83 25.29 20.69 16.59 14.19 12.25 9.75 7.54 5.77 4.86 4.32 3.50 3.06
Ours ($\rho=2.0$) 51.87 32.27 28.98 23.09 19.82 16.28 14.12 12.08 9.57 7.51 5.70 4.79 4.29 3.46 3.03
UniPC Uniform-$\lambda$ 92.24 68.91 45.57 33.01 24.71 19.12 15.23 12.42 9.03 6.35 4.56 3.80 3.46 3.01 2.85
Xue’24 49.77 36.07 21.33 14.94 12.39 12.21 14.55 9.51 7.98 6.28 4.55 3.80 3.47 3.01 2.85
Ours ($\rho=1.5$) 50.91 38.36 28.49 19.75 14.77 11.14 9.00 7.55 5.76 4.45 3.56 3.19 3.04 2.86 2.81
Ours ($\rho=2.0$) 43.55 28.43 22.24 17.02 13.48 10.70 8.89 7.46 5.76 4.51 3.56 3.17 3.01 2.82 2.80
To apply (11), we model the source distribution using a synthetic power-law covariance spectrum, which mimics the heavy-tailed eigenvalue decay characteristic of real image data [21]. Figure 1 presents the quantitative evaluation of the closed-form KL divergence (11) across varying step budgets $N$ (details in Appendix E). This comparison covers both VP and VE settings as considered in [20] with the same time discretization strategy given in (30). The results demonstrate that in both settings, the discretization strategy (30) with an appropriate choice of $\rho$ can achieve notably faster convergence compared to the widely used uniform-$\lambda$ baseline, significantly reducing the reverse sampling mismatch under limited budgets. These preliminary observations motivate further experimentation on real-world image generation tasks, as presented in the next section.
IV Experiments
Focusing on the practical image generation tasks using pretrained DMs, we evaluate the performance of our selected time discretization strategies in conjunction with off-the-shelf samplers. Experiments are conducted on CIFAR-10 and FFHQ-64 datasets using official model checkpoints from both EDM [6] and VP-SDE [20] architectures. We integrate our time discretization strategies with representative ODE samplers, specifically DPM-Solver++ [13] and UniPC [27], whose first-order update formulas correspond exactly to (3), and report the Fréchet Inception Distance (FID) score [4], a universally adopted metric in image generation tasks that quantifies the Wasserstein distance between the feature distributions of real images and 50k generated samples.
Comparison Schemes. We benchmark against two representative strategies with their officially available implementations: (i) Uniform-$\lambda$ [12], the widely adopted heuristic default for ODE samplers; (ii) Xue’24 [25], a state-of-the-art approach that attempts to balance between optimization overhead and generation performance for searching optimal steps. (iii) Ours, the power-uniform schemes (30) with $\rho=1.5$ and $2.0$.
The quantitative results are summarized in Table I. As shown, our approach consistently outperforms the default baseline by a significant margin, particularly under limited NFE budgets. Furthermore, in the majority of cases, it surpasses the competing search-based method. Notably, while the search-based method occasionally achieves competitive results in few-step regimes, such an edge diminishes significantly as the number of steps increases. Conversely, our approach maintains consistent superiority across all budgets without incurring any costly overhead, validating its effectiveness across varying model architectures and datasets.
V Conclusion
This work has presented a first-principles analysis of diffusion sampling, establishing an explicit connection between the noise schedule and sampling accuracy. Furthermore, it has demonstrated that the optimal schedule can be derived analytically for the Gaussian case, exhibiting a tangent law whose coefficient is determined by the eigenvalues of the source covariance matrix. Bridging theory and practice, we have shown the practical merit of our theoretical framework through relevant applications, illustrating how it serves as a useful tool for efficiently identifying improved time discretization strategies to enhance generation quality, compared to existing baselines. Of interest for future work in this area is to extend our theoretical framework to support guidance sampling and high-resolution generation tasks.
Appendix A Proof of Lemma 1
This proof shows that, under the optimal estimator given by (5) and (7), the explicit closed-form evolution trajectory of the distributions across reverse sampling steps can be derived.
Firstly, substituting (5) into the deterministic reverse sampling process (3) yields
$\hat{\mathbf{x}}_{t_{j-1}}$ $=\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\hat{\mathbf{x}}_{t_{j}}+\left(\sigma_{t_{j-1}}-\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\sigma_{t_{j}}\right)\frac{\hat{\mathbf{x}}_{t_{j}}-\alpha_{t_{j}}\mathbb{E}[\mathbf{x}_{0}|\hat{\mathbf{x}}_{t_{j}}]}{\sigma_{t_{i}}}$ (31)
$=\frac{\sigma_{t_{j-1}}}{\sigma_{t_{j}}}\hat{\mathbf{x}}_{t_{j}}+\left(\alpha_{t_{j-1}}-\frac{\sigma_{t_{j-1}}}{\sigma_{t_{j}}}\alpha_{t_{j}}\right)\mathbb{E}[\mathbf{x}_{0}|\hat{\mathbf{x}}_{t_{j}}].$
Under the Gaussian setup, the optimal posterior estimator (7) gives
$\mathbb{E}[\mathbf{x}_{0}|\hat{\mathbf{x}}_{t_{j}}]=\alpha_{t_{j}}\bm{\Sigma}_{\mathbf{x}}\left(\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t_{j}}^{2}\mathbf{I}\right)^{-1}\hat{\mathbf{x}}_{t_{j}}.$ (32)
Plugging (32) into (31) yields the linear update
$\hat{\mathbf{x}}_{t_{j-1}}$ $=\left(\frac{\sigma_{t_{j-1}}}{\sigma_{t_{j}}}\mathbf{I}+\left(\alpha_{t_{j-1}}-\frac{\sigma_{t_{j-1}}}{\sigma_{t_{j}}}\alpha_{t_{j}}\right)\alpha_{t_{j}}\bm{\Sigma}_{\mathbf{x}}\left(\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t_{j}}^{2}\mathbf{I}\right)^{-1}\right)\hat{\mathbf{x}}_{t_{j}}$
$=\left(\frac{\sigma_{t_{j-1}}}{\sigma_{t_{j}}}\left(\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t_{j}}^{2}\mathbf{I}\right)+\left(\alpha_{t_{j-1}}\alpha_{t_{j}}\bm{\Sigma}_{\mathbf{x}}-\frac{\sigma_{t_{j-1}}}{\sigma_{t_{j}}}\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}\right)\right)\left(\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t_{j}}^{2}\mathbf{I}\right)^{-1}\hat{\mathbf{x}}_{t_{j}}$
$=\left(\sigma_{t_{j-1}}\sigma_{t_{j}}\mathbf{I}+\alpha_{t_{j-1}}\alpha_{t_{j}}\bm{\Sigma}_{\mathbf{x}}\right)\left(\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t_{j}}^{2}\mathbf{I}\right)^{-1}\hat{\mathbf{x}}_{t_{j}}.$ (33)
Since $\bm{\Sigma}_{\mathbf{x}}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{-1}$ with $\bm{\Lambda}=\mathrm{diag}(\mu_{1},\ldots,\mu_{k})$, we have $(\alpha_{t_{j}}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t_{j}}^{2}\mathbf{I})^{-1}=\mathbf{U}(\alpha_{t_{j}}^{2}\bm{\Lambda}+\sigma_{t_{j}}^{2}\mathbf{I})^{-1}\mathbf{U}^{-1}$. Then the updata formula (33) becomes
$\hat{\mathbf{x}}_{t_{j-1}}$ $=\mathbf{U}\left(\sigma_{t_{j-1}}\sigma_{t_{j}}\mathbf{I}+\alpha_{t_{j-1}}\alpha_{t_{j}}\bm{\Lambda}\right)\mathbf{U}^{-1}\mathbf{U}(\alpha_{t_{j}}^{2}\bm{\Lambda}+\sigma_{t_{j}}^{2}\mathbf{I})^{-1}\mathbf{U}^{-1}\hat{\mathbf{x}}_{t_{j}}$
$=\mathbf{U}\left(\sigma_{t_{j-1}}\sigma_{t_{j}}\mathbf{I}+\alpha_{t_{j-1}}\alpha_{t_{j}}\bm{\Lambda}\right)(\alpha_{t_{j}}^{2}\bm{\Lambda}+\sigma_{t_{j}}^{2}\mathbf{I})^{-1}\mathbf{U}^{-1}\hat{\mathbf{x}}_{t_{j}}$
$\triangleq\mathbf{U}\mathbf{D}_{j}\mathbf{U}^{-1}\hat{\mathbf{x}}_{t_{j}},$ (34)
where the diagonal matrix $\mathbf{D}_{j}$ is
$\mathbf{D}_{j}=\left(\alpha_{t_{j-1}}\alpha_{t_{j}}\bm{\Lambda}+\sigma_{t_{j-1}}\sigma_{t_{j}}\mathbf{I}\right)\left(\alpha_{t_{j}}^{2}\bm{\Lambda}+\sigma_{t_{j}}^{2}\mathbf{I}\right)^{-1}=\mathrm{diag}(d_{j,1},\ldots,d_{j,k}),$
with the diagonal entries
$d_{j,\ell}=\frac{\alpha_{t_{j-1}}\alpha_{t_{j}}\mu_{\ell}+\sigma_{t_{j-1}}\sigma_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}},\qquad\ell=1,\ldots,k.$ (35)
Then, using the diagonalized linear update (A) repeatedly with the time discretization sequence $\{t_{i}\}_{i=0}^{N}$ yields
$\hat{\mathbf{x}}_{t_{0}}=\mathbf{U}\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)\mathbf{U}^{-1}\hat{\mathbf{x}}_{t_{N}},$ (36)
By (6), in the practical reverse sampling, the reverse process is initialized from the terminal prior:
$\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{t_{N}}^{2}\mathbf{I}),\quad\text{where}\quad\sigma_{t_{N}}=\begin{cases}1&\text{VP setting},\\ \sigma_{\max}&\text{VE setting}.\end{cases}$ (37)
Since the update (36) is affine transformation on a Gaussian distribution, $\hat{\mathbf{x}}_{t_{0}}$ remains Gaussian:
$p(\hat{\mathbf{x}}_{t_{0}})=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),$ (38)
where the diagonal matrix $\mathbf{M}$ is
$\mathbf{M}$ $\triangleq\sigma_{t_{N}}^{2}\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)^{\!\top}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (39)
with its diagonal entries
$m_{\ell}=\sigma_{t_{N}}^{2}\prod_{j=1}^{N}d_{j,\ell}^{2},\qquad\ell=1,\ldots,k,$ (40)
where $d_{j,\ell}$ is given in (35). This yields the result stated in Lemma 1.
The KL divergence between two zero-mean Gaussians is
$D_{\mathrm{KL}}(\mathcal{N}(0,\Sigma_{p})\|\mathcal{N}(0,\Sigma_{q}))=\frac{1}{2}\left(\mathrm{tr}(\mathbf{S})-\log\det(\mathbf{S})-k\right),$
where $\mathbf{S}=\Sigma_{q}^{-1}\Sigma_{p}$. Finally, combining the distribution $q(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$ with $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$, we can immediately obtain
$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)$ $=D_{\mathrm{KL}}\!\left(\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right)\,\|\,\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{N}\mathbf{U}^{-1}\right)\right)$
$=\frac{1}{2}\left(\mathrm{tr}(\mathbf{N}^{-1}\mathbf{M})-\log\det(\mathbf{N}^{-1}\mathbf{M})-k\right)$
$=\frac{1}{2}\sum_{\ell=1}^{k}\left(\frac{m_{\ell}}{n_{\ell}}-\log\frac{m_{\ell}}{n_{\ell}}-1\right),$ (41)
where $m_{\ell}$ and $n_{\ell}$ are respectively given in (40) and (10).
Appendix B Proof of Lemma 2
For a fixed mode $\ell\in\{1,\ldots,k\}$, recall
$S_{\ell}^{N}=\sum_{j=1}^{N}\log\!\left(\frac{\alpha_{t_{j-1}}\alpha_{t_{j}}\mu_{\ell}+\sigma_{t_{j-1}}\sigma_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\right).$ (42)
Proposition 1.
Let $t_{j}\triangleq j/N$ be a uniform grid on $[0,1]$ with step size $h\triangleq 1/N$. Assume that $\alpha,\sigma$ are sufficiently smooth. Then, the sum $S_{\ell}^{N}$ admits the following expansion as $h\to 0$:
$S_{\ell}^{N}=\sum_{j=1}^{N}\Bigl(F_{\ell}(t_{j})\,h+G_{\ell}(t_{j})\,h^{2}\Bigr)+\mathcal{O}(h^{2}),$ (43)
where the coefficients are given by
$F_{\ell}(t_{j})$ $=-\frac{\alpha_{t_{j}}\dot{\alpha}_{t_{j}}\mu_{\ell}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}}{\alpha^{2}_{t_{j}}\mu_{\ell}+\sigma^{2}_{t_{j}}},$ (44)
$G_{\ell}(t_{j})$ $=\frac{\alpha_{t_{j}}\ddot{\alpha}_{t_{j}}\mu_{\ell}+\sigma_{t_{j}}\ddot{\sigma}_{t_{j}}}{2\bigl(\alpha^{2}_{t_{j}}\mu_{\ell}+\sigma^{2}_{t_{j}}\bigr)}-\frac{\bigl(\alpha_{t_{j}}\dot{\alpha}_{t_{j}}\mu_{\ell}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}\bigr)^{2}}{2\bigl(\alpha^{2}_{t_{j}}\mu_{\ell}+\sigma^{2}_{t_{j}}\bigr)^{2}}.$ (45)
Proof.
Using the Taylor expansion at $t_{j}$ (with $t_{j-1}=t_{j}-h$), we have
$\alpha_{t_{j-1}}$ $=\alpha(t_{j}-h)=\alpha_{t_{j}}-h\,\dot{\alpha}_{t_{j}}+\frac{h^{2}}{2}\,\ddot{\alpha}_{t_{j}}+\mathcal{O}(h^{3}),$
$\sigma_{t_{j-1}}$ $=\sigma(t_{j}-h)=\sigma_{t_{j}}-h\,\dot{\sigma}_{t_{j}}+\frac{h^{2}}{2}\,\ddot{\sigma}_{t_{j}}+\mathcal{O}(h^{3}),$ (46)
where $\dot{\alpha}_{t_{j}}\triangleq(\mathrm{d}\alpha/\mathrm{d}t)(t_{j})$ and $\ddot{\alpha}_{t_{j}}\triangleq(\mathrm{d}^{2}\alpha/\mathrm{d}t^{2})(t_{j})$, and similarly for $\sigma$.
Substituting (B) into the numerator term of the summand in (42), we obtain
$\alpha_{t_{j-1}}\alpha_{t_{j}}\mu_{\ell}+\sigma_{t_{j-1}}\sigma_{t_{j}}$ $=\left(\alpha_{t_{j}}-h\,\dot{\alpha}_{t_{j}}+\frac{h^{2}}{2}\,\ddot{\alpha}_{t_{j}}\right)\alpha_{t_{j}}\mu_{\ell}+\left(\sigma_{t_{j}}-h\,\dot{\sigma}_{t_{j}}+\frac{h^{2}}{2}\,\ddot{\sigma}_{t_{j}}\right)\sigma_{t_{j}}+\mathcal{O}(h^{3})$
$=\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}-h\bigl(\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}\bigr)+\frac{h^{2}}{2}\bigl(\mu_{\ell}\alpha_{t_{j}}\ddot{\alpha}_{t_{j}}+\sigma_{t_{j}}\ddot{\sigma}_{t_{j}}\bigr)+\mathcal{O}(h^{3}).$ (47)
Then, using (B), the sum $S_{\ell}^{N}$ becomes
$S_{\ell}^{N}$ $=\sum_{j=1}^{N}\log\left(\frac{\alpha_{t_{j-1}}\alpha_{t_{j}}\mu_{\ell}+\sigma_{t_{j-1}}\sigma_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\right)$
$=\sum_{j=1}^{N}\log\left(1+\frac{-h\bigl(\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}\bigr)+\frac{h^{2}}{2}\bigl(\mu_{\ell}\alpha_{t_{j}}\ddot{\alpha}_{t_{j}}+\sigma_{t_{j}}\ddot{\sigma}_{t_{j}}\bigr)}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}+\mathcal{O}(h^{3})\right)$
$\triangleq\sum_{j=1}^{N}\log\left(1+\delta_{j}\right),$ (48)
where the term $\delta_{j}$ is expanded as
$\delta_{j}$ $=\frac{-h\bigl(\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}\bigr)+\frac{h^{2}}{2}\bigl(\mu_{\ell}\alpha_{t_{j}}\ddot{\alpha}_{t_{j}}+\sigma_{t_{j}}\ddot{\sigma}_{t_{j}}\bigr)}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}+\mathcal{O}(h^{3})$
$=-\frac{\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\,h+\frac{\mu_{\ell}\alpha_{t_{j}}\ddot{\alpha}_{t_{j}}+\sigma_{t_{j}}\ddot{\sigma}_{t_{j}}}{2(\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2})}\,h^{2}+\mathcal{O}(h^{3})$
$\triangleq a_{j}h+b_{j}h^{2}+\mathcal{O}(h^{3}).$ (49)
Since $\alpha(\cdot)$ and $\sigma(\cdot)$ are smooth, we have $\delta_{j}=\mathcal{O}(h)$ as $h\to 0$. Consequently, squaring this expression yields
$\delta_{j}^{2}=a_{j}^{2}h^{2}+\mathcal{O}(h^{3})=\left(\frac{\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\right)^{2}h^{2}+\mathcal{O}(h^{3}).$ (50)
Applying the Taylor expansion $\log(1+x)=x-\frac{x^{2}}{2}+\mathcal{O}(x^{3})$, we get
$\log(1+\delta_{j})$ $=\delta_{j}-\frac{1}{2}\delta_{j}^{2}+\mathcal{O}(\delta_{j}^{3})=\left(a_{j}h+b_{j}h^{2}\right)-\frac{1}{2}\left(a_{j}^{2}h^{2}\right)+\mathcal{O}(h^{3})$
$=-\frac{\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}}{\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2}}\,h+\left[\frac{\mu_{\ell}\alpha_{t_{j}}\ddot{\alpha}_{t_{j}}+\sigma_{t_{j}}\ddot{\sigma}_{t_{j}}}{2(\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2})}-\frac{\bigl(\mu_{\ell}\alpha_{t_{j}}\dot{\alpha}_{t_{j}}+\sigma_{t_{j}}\dot{\sigma}_{t_{j}}\bigr)^{2}}{2(\alpha_{t_{j}}^{2}\mu_{\ell}+\sigma_{t_{j}}^{2})^{2}}\right]h^{2}+\mathcal{O}(h^{3})$
$\triangleq F_{\ell}(t_{j})h+G_{\ell}(t_{j})h^{2}+\mathcal{O}(h^{3}),$ (51)
where we used the fact that $\mathcal{O}(\delta_{j}^{3})=\mathcal{O}(h^{3})$. Finally, substituting (B) into $S_{\ell}^{N}=\sum_{j=1}^{N}\log(1+\delta_{j})$, and noting that the summation of local $\mathcal{O}(h^{3})$ error terms over $N\propto h^{-1}$ steps results in a global error of $\mathcal{O}(h^{2})$, we obtain
$S_{\ell}^{N}=\sum_{j=1}^{N}\Bigl(F_{\ell}(t_{j})\,h+G_{\ell}(t_{j})\,h^{2}\Bigr)+\mathcal{O}(h^{2}),$ (52)
which concludes the proof. ∎
To proceed with the derivation, we first introduce the Euler–Maclaurin formula [30]. Specifically, a sum of the form $\sum_{j=1}^{N}f(jh)$ with $h=1/N$ admits the following asymptotic expansion:
$\sum_{j=1}^{N}f(jh)\,h=\int_{0}^{1}f(s)\,\mathrm{d}s+\frac{h}{2}\bigl(f(1)-f(0)\bigr)+\frac{h^{2}}{12}\bigl(f^{\prime}(1)-f^{\prime}(0)\bigr)+R(f,h),$ (53)
where the residual term satisfies $R(f,h)=\mathcal{O}(h^{3})$ as $h\to 0$.
Applying (53) to the Riemann sums in Proposition 1 yields, as $h\to 0$,
$\sum_{j=1}^{N}F_{\ell}(t_{j})\,h$ $=\int_{0}^{1}F_{\ell}(t)\,\mathrm{d}t+\frac{h}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)+\mathcal{O}(h^{2}),$ (54)
$\sum_{j=1}^{N}G_{\ell}(t_{j})\,h$ $=\int_{0}^{1}G_{\ell}(t)\,\mathrm{d}t+\mathcal{O}(h),$
where the function $F_{\ell}(t)$ and $G_{\ell}(t)$ are
$F_{\ell}(t)$ $=-\frac{\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma}}{\alpha^{2}\mu_{\ell}+\sigma^{2}},$ (55)
$G_{\ell}(t)$ $=\frac{\alpha\ddot{\alpha}\mu_{\ell}+\sigma\ddot{\sigma}}{2\left(\alpha^{2}\mu_{\ell}+\sigma^{2}\right)}-\frac{\left(\alpha\dot{\alpha}\mu_{\ell}+\sigma_{t_{j}}\dot{\sigma}\right)^{2}}{2\left(\alpha^{2}\mu_{\ell}+\sigma^{2}\right)^{2}}.$ (56)
Recalling the expansion in (43) and substituting the approximations from (54), we obtain
$S_{\ell}^{N}$ $=\sum_{j=1}^{N}\Bigl(F_{\ell}(t_{j})\,h+G_{\ell}(t_{j})\,h^{2}\Bigr)+\mathcal{O}(h^{2})$
$=\sum_{j=1}^{N}F_{\ell}(t_{j})\,h+h\sum_{j=1}^{N}G_{\ell}(t_{j})\,h+\mathcal{O}(h^{2})$
$=\int_{0}^{1}F_{\ell}(t)\,\mathrm{d}t+\frac{h}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)+h\left(\int_{0}^{1}G_{\ell}(t)\,\mathrm{d}t+\mathcal{O}(h)\right)+\mathcal{O}(h^{2})$
$=\int_{0}^{1}F_{\ell}(t)\,\mathrm{d}t+E_{\ell}^{1}\,h+\mathcal{O}(h^{2}),$ (57)
where the leading order coefficient is given by
$E_{\ell}^{1}=\frac{1}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)+\int_{0}^{1}G_{\ell}(t)\,\mathrm{d}t.$ (58)
Define the continuous counterpart
$I_{\ell}\triangleq\int_{0}^{1}F_{\ell}(t)\,\mathrm{d}t.$ (59)
Then by (57) and (59), the discretization error $r_{\ell}\triangleq S_{\ell}^{N}-I_{\ell}$ satisfies
$r_{\ell}=\frac{E_{\ell}^{1}}{N}+\mathcal{O}\!\left(\frac{1}{N^{2}}\right).$ (60)
To explicitly quantify the error, we proceed to derive the analytical forms for the integral term $I_{\ell}$ and the leading-order coefficient $E_{\ell}^{1}$.
B-A Closed Form of Integral Term $I_{\ell}$
Substituting the definition of $F_{\ell}(t)$ (55) and using the differential relation $\mathrm{d}(\alpha^{2}\mu_{\ell}+\sigma^{2})=2(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})\,\mathrm{d}t$, the integral admits an analytic closed-form solution:
$I_{\ell}$ $=\int_{0}^{1}F_{\ell}(t)\,\mathrm{d}t$
$=\int_{0}^{1}-\frac{\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma}}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\,\mathrm{d}t$
$=-\frac{1}{2}\int_{0}^{1}\frac{\mathrm{d}(\alpha^{2}\mu_{\ell}+\sigma^{2})}{\alpha^{2}\mu_{\ell}+\sigma^{2}}$
$=-\frac{1}{2}\log(\alpha^{2}\mu_{\ell}+\sigma^{2})\Big|_{t=0}^{t=1}$
$=\frac{1}{2}\log\!\left(\frac{\alpha^{2}(0)\mu_{\ell}+\sigma^{2}(0)}{\alpha^{2}(1)\mu_{\ell}+\sigma^{2}(1)}\right).$ (61)
To interpret this result, we first recall that $n_{\ell}\triangleq\alpha^{2}(0)\mu_{\ell}+\sigma^{2}(0)$ (10). We now consider the following two boundary conditions, as we mentioned before in Section II-A.
  • Variance Preserving (VP): The boundary conditions are specified as $\alpha(0)=1,\sigma(0)=0$ and $\alpha(1)=0,\sigma(1)=1$. Substituting these into (61) yields
    $I_{\ell}\big|_{\text{VP}}=\frac{1}{2}\log\left(\frac{n_{\ell}}{0\cdot\mu_{\ell}+1}\right)=\frac{1}{2}\log n_{\ell}.$ (62)
  • Variance Exploding (VE): The boundary conditions are $\alpha(t)\equiv 1$, $\sigma(0)=0$, and $\sigma(1)=\sigma_{\max}$. In this case, the integral (61) becomes
    $I_{\ell}\big|_{\text{VE}}=\frac{1}{2}\log\!\left(\frac{n_{\ell}}{\mu_{\ell}+\sigma_{\max}^{2}}\right).$ (63)
    Consider the regime where $\sigma_{\max}^{2}\gg\mu_{\ell}$ (i.e., $\mu_{\ell}/\sigma_{\max}^{2}\to 0$). By applying the Taylor expansion $\log(A+x)=\log A+x/A+\mathcal{O}(x^{2}/A^{2})$, we obtain the asymptotic form:
    $I_{\ell}\big|_{\text{VE}}+\log\sigma_{\max}=\frac{1}{2}\log n_{\ell}+\mathcal{O}\!\left(\frac{\mu_{\ell}}{\sigma_{\max}^{2}}\right).$ (64)
Combining the result for VP (62) and the asymptotic approximation for VE (64), we establish the following unified relationship:
$\frac{1}{2}\log n_{\ell}\approx I_{\ell}+\log\sigma_{N}.$ (65)
where $\sigma_{N}$ is defined in (37). Note that the relation in (65) is exact for the VP setting (where $\sigma_{N}=1$), and holds asymptotically for the VE setting (where $\sigma_{N}=\sigma_{\max}$) in the high-noise setting ($\sigma_{\max}^{2}\gg\mu_{\ell}$).
Recalling the discrete counterpart from (40) and (42), we have the exact identity
$\frac{1}{2}\log m_{\ell}=S_{\ell}^{N}+\log\sigma_{N}.$ (66)
Subtracting (65) from (66), we obtain
$\frac{1}{2}\log\frac{m_{\ell}}{n_{\ell}}\approx S_{\ell}^{N}-I_{\ell}.$
Consequently, to minimize the KL divergence given in (A), which is determined by the discrepancy between $m_{\ell}$ and $n_{\ell}$, effectively reduces to minimizing the discretization error between the Riemann sum $S_{\ell}^{N}$ and the integral $I_{\ell}$, defined as $r_{\ell}$ in (60).
B-B Simplification of Leading Order Coefficient $E_{\ell}^{1}$
We decompose the integrand $G_{\ell}(t)$ defined in (56) into two parts, denoted as $A_{\ell}(t)$ and $B_{\ell}(t)$:
$G_{\ell}(t)=\underbrace{\frac{\alpha\ddot{\alpha}\mu_{\ell}+\sigma\ddot{\sigma}}{2(\alpha^{2}\mu_{\ell}+\sigma^{2})}}_{A_{\ell}(t)}\;-\;\underbrace{\frac{(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})^{2}}{2(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}}_{B_{\ell}(t)}.$ (67)
Substituting this decomposition into the expression for $E_{\ell}^{1}$ (58), we obtain
$E_{\ell}^{1}$ $=\frac{1}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)+\int_{0}^{1}G_{\ell}(t)\,\mathrm{d}t$
$=\frac{1}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)+\int_{0}^{1}A_{\ell}(t)\,\mathrm{d}t-\int_{0}^{1}B_{\ell}(t)\,\mathrm{d}t.$ (68)
Applying integration by parts to $\int A_{\ell}(t)\,\mathrm{d}t$ yields
$\int_{0}^{1}A_{\ell}(t)\,\mathrm{d}t$ $=\int_{0}^{1}\frac{1}{2}\left(\frac{\alpha\mu_{\ell}}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\ddot{\alpha}+\frac{\sigma}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\ddot{\sigma}\right)\,\mathrm{d}t$
$=\frac{1}{2}\int_{0}^{1}\frac{\alpha\mu_{\ell}}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\,\mathrm{d}(\dot{\alpha})+\frac{1}{2}\int_{0}^{1}\frac{\sigma}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\,\mathrm{d}(\dot{\sigma})$
$=\frac{1}{2}\left[\frac{\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma}}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\right]_{0}^{1}-\frac{1}{2}\int_{0}^{1}\left(\dot{\alpha}\frac{\mathrm{d}}{\mathrm{d}t}\frac{\alpha\mu_{\ell}}{\alpha^{2}\mu_{\ell}+\sigma^{2}}+\dot{\sigma}\frac{\mathrm{d}}{\mathrm{d}t}\frac{\sigma}{\alpha^{2}\mu_{\ell}+\sigma^{2}}\right)\mathrm{d}t$
$=-\frac{1}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)-\frac{1}{2}\int_{0}^{1}\frac{(\dot{\alpha}^{2}\mu_{\ell}+\dot{\sigma}^{2})(\alpha^{2}\mu_{\ell}+\sigma^{2})-2(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\mathrm{d}t$
$=-\frac{1}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)-\frac{1}{2}\int_{0}^{1}\frac{(\dot{\alpha}^{2}\mu_{\ell}+\dot{\sigma}^{2})(\alpha^{2}\mu_{\ell}+\sigma^{2})-2(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t,$ (69)
where we identified the boundary term as $\frac{1}{2}(F_{\ell}(1)-F_{\ell}(0))$ based on the definition of $F_{\ell}(t)$ in (55).
Substituting the result from (B-B) back into the expression of $E_{\ell}^{1}$ given in (B-B), we obtain
$E_{\ell}^{1}$ $=\frac{1}{2}\bigl(F_{\ell}(1)-F_{\ell}(0)\bigr)+\int_{0}^{1}A_{\ell}(t)\,\mathrm{d}t-\int_{0}^{1}B_{\ell}(t)\,\mathrm{d}t$
$=-\frac{1}{2}\int_{0}^{1}\frac{(\dot{\alpha}^{2}\mu_{\ell}+\dot{\sigma}^{2})(\alpha^{2}\mu_{\ell}+\sigma^{2})-2(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t-\int_{0}^{1}\frac{(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})^{2}}{2(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t$
$=-\frac{1}{2}\int_{0}^{1}\frac{(\dot{\alpha}^{2}\mu_{\ell}+\dot{\sigma}^{2})(\alpha^{2}\mu_{\ell}+\sigma^{2})-(\alpha\dot{\alpha}\mu_{\ell}+\sigma\dot{\sigma})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t$
$=-\frac{\mu_{\ell}}{2}\int_{0}^{1}\frac{(\alpha\dot{\sigma}-\sigma\dot{\alpha})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t.$ (70)
This final expression confirms that the leading order coefficient $E_{\ell}^{1}$ is non-positive for arbitrary noise schedules $\alpha(t),\sigma(t)$, provided that $\mu_{\ell}>0$.
Appendix C Supplementary Details of Theorem 1
In this appendix, we extend the result of Theorem 1 to the VE setting. While the objective functional remains identical to that of the VP setting (21), the primary difference lies in the boundary condition at $t=1$.
Given the noise schedule $\alpha(t)\equiv 1$, $\sigma(0)=0$, and $\sigma(1)=\sigma_{\max}$ for the VE setting, along with the definition $\eta(t)=\sigma(t)/\alpha(t)$, we derive the boundary conditions as follows:
$\eta(0)=0\quad\text{and}\quad\eta(1)=\sigma_{\max}.$ (71)
Based on the definition $Q_{\ell}(t)\triangleq\frac{1}{\sqrt{\mu_{\ell}}}\arctan\!\left(\frac{\eta(t)}{\sqrt{\mu_{\ell}}}\right)$ from (22), the transformed boundary conditions become
$Q_{\ell}(0)=0\quad\text{and}\quad Q_{\ell}(1)=\frac{1}{\sqrt{\mu_{\ell}}}\arctan\!\left(\frac{\sigma_{\max}}{\sqrt{\mu_{\ell}}}\right).$ (72)
Recall that the Euler–Lagrange equation associated with the Lagrangian $L=\dot{Q}_{\ell}^{2}$ (24) dictates that the optimal trajectory is linear: $Q_{\ell}(t)=ct+d.$
By imposing the boundary conditions (72) on this linear solution, we obtain the explicit optimal trajectory in the transformed space:
$Q_{\ell}(t)=\frac{t}{\sqrt{\mu_{\ell}}}\arctan\!\left(\frac{\sigma_{\max}}{\sqrt{\mu_{\ell}}}\right).$ (73)
Subsequently, by substituting the optimal trajectory (73) into the definition (22), we derive the corresponding tangent law schedule for the VE setting $\eta_{\ell}(t)\big|_{\text{VE}}$:
$\eta_{\ell}(t)\big|_{\text{VE}}=\sqrt{\mu_{\ell}}\tan\!\left(t\cdot\arctan\!\left(\frac{\sigma_{\max}}{\sqrt{\mu_{\ell}}}\right)\right),$ (74)
so when $\sigma_{\max}^{2}\gg\mu_{\ell}$, we have $\arctan\!\left(\frac{\sigma_{\max}}{\sqrt{\mu_{\ell}}}\right)\approx\frac{\pi}{2}$, implying that $\eta_{\ell}(t)\big|_{\text{VE}}$ (74) approaches the VP solution $\eta_{\ell}(t)\big|_{\text{VP}}$ (20) given in Definition 1.
Appendix D Supplementary Details of Theorem 2
In this appendix, we derive an explicit expression for the leading-order coefficient $E_{\ell}^{1}$ under the parameterized tangent law schedule. Additionally, we provide supplementary numerical results regarding Theorem 2.
D-A Derivation of Leading-Order Coefficient
Building upon the tangent laws derived for VP (20) and VE (74) in Theorem 1, we formulate a unified parameterized tangent law schedule applicable to both settings:
$\eta_{\gamma}(t)\triangleq\frac{\sigma_{\gamma}(t)}{\alpha_{\gamma}(t)}=\sqrt{\gamma}\,\tan\!\bigl(\Theta_{\gamma}t\bigr),\quad\gamma>0,$ (75)
where the factor $\Theta_{\gamma}$ is determined by the endpoint condition $\eta_{\gamma}(1)=\eta_{1}$:
$\Theta_{\gamma}\triangleq\arctan\!\left(\frac{\eta_{1}}{\sqrt{\gamma}}\right).$ (76)
We specify the boundary values $\eta_{1}$ for the following two settings:
  • Variance Preserving (VP): Since $\alpha(1)=0$ and $\sigma(1)=1$, we have $\eta_{1}\to\infty$. Consequently, the angle becomes
    $\Theta_{\gamma}\big|_{\text{VP}}=\frac{\pi}{2}.$
  • Variance Exploding (VE): Since $\alpha(t)\equiv 1$ and $\sigma(1)=\sigma_{\max}$, we have $\eta_{1}=\sigma_{\max}$,
    $\Theta_{\gamma}\big|_{\text{VE}}=\arctan\!\left(\frac{\sigma_{\max}}{\sqrt{\gamma}}\right)=\frac{\pi}{2}-\arctan\!\left(\frac{\sqrt{\gamma}}{\sigma_{\max}}\right).$
Same as we mentioned before in Appendix B-A and Appendix C, under the condition of $\sigma_{\max}^{2}\gg\gamma$, the VP setting solution can be regarded as an accurate approximation of the VE setting.
Substituting the parameterized tangent law schedule $\eta_{\gamma}(t)=\sqrt{\gamma}\tan(\Theta_{\gamma}t)$ into the expression of $E_{\ell}^{1}$ given in (70), we proceed as follows:
$E_{\ell}^{1}(\gamma)$ $=-\frac{\mu_{\ell}}{2}\int_{0}^{1}\frac{(\alpha\dot{\sigma}-\sigma\dot{\alpha})^{2}}{(\alpha^{2}\mu_{\ell}+\sigma^{2})^{2}}\,\mathrm{d}t$
$=-\frac{\mu_{\ell}}{2}\int_{0}^{1}\frac{\dot{\eta}(t)^{2}}{(\mu_{\ell}+\eta(t)^{2})^{2}}\,\mathrm{d}t$
$=-\frac{\mu_{\ell}}{2}\int_{0}^{1}\frac{\left[\frac{\mathrm{d}}{\mathrm{d}t}\left(\sqrt{\gamma}\tan\left(\Theta_{\gamma}t\right)\right)\right]^{2}}{\left(\mu_{\ell}+\gamma\tan^{2}\left(\Theta_{\gamma}t\right)\right)^{2}}\,\mathrm{d}t$
$=-\frac{\mu_{\ell}}{2}\int_{0}^{1}\frac{\gamma\Theta_{\gamma}^{2}\sec^{4}(\Theta_{\gamma}t)}{\left(\frac{\mu_{\ell}\cos^{2}(\Theta_{\gamma}t)+\gamma\sin^{2}(\Theta_{\gamma}t)}{\cos^{2}(\Theta_{\gamma}t)}\right)^{2}}\,\mathrm{d}t$
$=-\frac{\mu_{\ell}\gamma\Theta_{\gamma}^{2}}{2}\int_{0}^{1}\frac{\sec^{4}(\Theta_{\gamma}t)\cdot\cos^{4}(\Theta_{\gamma}t)}{\bigl(\mu_{\ell}\cos^{2}(\Theta_{\gamma}t)+\gamma\sin^{2}(\Theta_{\gamma}t)\bigr)^{2}}\,\mathrm{d}t$
$=-\frac{\mu_{\ell}\gamma\Theta_{\gamma}}{2}\int_{0}^{\Theta_{\gamma}}\frac{1}{\bigl(\mu_{\ell}\cos^{2}x+\gamma\sin^{2}x\bigr)^{2}}\,\mathrm{d}x$
$=-\frac{\mu_{\ell}\gamma\Theta_{\gamma}}{2}\int_{0}^{\Theta_{\gamma}}\frac{1}{\bigl(\mu_{\ell}+(\gamma-\mu_{\ell})\sin^{2}x\bigr)^{2}}\,\mathrm{d}x.$ (77)
The standard integral result from [2, formula 2.563.1 and 2.562.1 in pp.177] is
$\int\frac{\mathrm{d}x}{(a+b\sin^{2}x)^{2}}$ $=\frac{1}{2a(a+b)}\left[(2a+b)\int\frac{\mathrm{d}x}{a+b\sin^{2}x}+\frac{b\sin x\cos x}{a+b\sin^{2}x}\right]$
$=\frac{1}{2a(a+b)}\left[\frac{(2a+b)\operatorname{sign}(a)}{\sqrt{a(a+b)}}\arctan\left(\sqrt{\frac{a+b}{a}}\tan x\right)+\frac{b\sin x\cos x}{a+b\sin^{2}x}\right](b>-a),$ (78)
where $\operatorname{sign}$ denotes the sign function. Applying this to $E_{\ell}^{1}(\gamma)$ (77) with $a=\mu_{\ell}$ and $b=\gamma-\mu_{\ell}$, we obtain
$E_{\ell}^{1}(\gamma)$ $=-\frac{\mu_{\ell}\gamma\Theta_{\gamma}}{2}\int_{0}^{\Theta_{\gamma}}\frac{1}{\bigl(\mu_{\ell}+(\gamma-\mu_{\ell})\sin^{2}x\bigr)^{2}}\,\mathrm{d}x$
$=-\frac{\mu_{\ell}\gamma\Theta_{\gamma}}{2}\cdot\frac{1}{2\mu_{\ell}\gamma}\left[\frac{\mu_{\ell}+\gamma}{\sqrt{\mu_{\ell}\gamma}}\arctan\left(\sqrt{\frac{\gamma}{\mu_{\ell}}}\tan\Theta_{\gamma}\right)+\frac{(\gamma-\mu_{\ell})\sin\Theta_{\gamma}\cos\Theta_{\gamma}}{\mu_{\ell}\cos^{2}\Theta_{\gamma}+\gamma\sin^{2}\Theta_{\gamma}}\right]$
$=-\frac{\Theta_{\gamma}}{4}\left[\frac{\mu_{\ell}+\gamma}{\sqrt{\mu_{\ell}\gamma}}\arctan\left(\sqrt{\frac{\gamma}{\mu_{\ell}}}\tan\Theta_{\gamma}\right)+\frac{(\gamma-\mu_{\ell})\tan\Theta_{\gamma}}{\mu_{\ell}+\gamma\tan^{2}\Theta_{\gamma}}\right].$ (79)
Finally, in the VP setting where $\Theta_{\gamma}=\frac{\pi}{2}$, we observe the following limiting behaviors
$\arctan\!\left(\sqrt{\frac{\gamma}{\mu_{\ell}}}\tan\Theta_{\gamma}\right)\to\frac{\pi}{2},\qquad\frac{(\gamma-\mu_{\ell})\tan\Theta_{\gamma}}{\mu_{\ell}+\gamma\tan^{2}\Theta_{\gamma}}\to 0.$
Substituting these limits yields the explicit expression
$E_{\ell}^{1}(\gamma)\big|_{\text{VP}}=-\frac{\pi/2}{4}\left(\frac{\mu_{\ell}+\gamma}{\sqrt{\mu_{\ell}\gamma}}\cdot\frac{\pi}{2}\right)=-\frac{\pi^{2}}{16}\left(\sqrt{\frac{\mu_{\ell}}{\gamma}}+\sqrt{\frac{\gamma}{\mu_{\ell}}}\right),$ (80)
thereby recovering the expression presented in (27).
D-B Numerical Results on Theorem 2
To assess the performance of the $\gamma^{\star}$ parameterized tangent law noise schedule derived in Theorem 2, we conduct a two-fold evaluation based on the closed-form KL divergence (11), with results summarized in Figure 2. First, we compare our parameterized tangent law schedule (using $\gamma^{\star}$) against representative noise schedules widely employed in DMs (Figure 2 top row). The results indicate that our method achieves faster convergence in both low-dimensional and high-dimensional synthetic Gaussian sources (details in Appendix E). Second, we perform an ablation study on the hyperparameter $\gamma$ (Figure 2 bottom row). The empirical results align with our theoretical analysis, showing that the derived $\gamma^{\star}$ (26) leads to the best performance compared to other choices.
(a) Low-dim ($k=128$).
(b) High-dim ($k=1280$, log scale).
(c) Low-dim ($k=128$, log scale).
(d) High-dim ($k=1280$, log scale).
Figure 2: Performance analysis of the parameterized tangent law schedule in Gaussian settings. We report the KL divergence versus the number of time discretization steps NN. Top Row: Comparison of our parameterized tangent law schedule (using $\gamma^{\star}$) against existing noise schedules [5, 20, 11, 14]. Bottom Row: Ablation study of different $\gamma$ values, verifying the optimality of our derived $\gamma^{\star}$. The left and right columns correspond to low-dimensional ($k=128$) and high-dimensional ($k=1280$) cases, respectively. The results demonstrate the effectiveness of our derived $\gamma^{\star}$ in Theorem 2.
Appendix E Implementation Details
In this appendix, we provide implementation details for the experiments discussed in Section III-C and Section IV, along with more qualitative results and visual samples.
Time Discretization Strategy
The power-uniform discretization strategy is formulated as:
$\lambda_{i}^{(\rho)}=\left(\lambda_{t_{N}}^{1/\rho}+\frac{i}{N}\left(\lambda_{t_{0}}^{1/\rho}-\lambda_{t_{N}}^{1/\rho}\right)\right)^{\rho},\quad i=0,\dots,N,$ (81)
where $\lambda(t)=\log(\alpha(t)/\sigma(t))$ represents the half-logSNR. The boundary values $\lambda_{t_{0}}$ and $\lambda_{t_{N}}$ correspond to the start and end points of the sampling trajectory, determined by the noise schedules employed during training (4).
To accommodate negative values (specifically when $\lambda<0$), we adopt the sign-preserving power, defined as:
$x^{\rho}\triangleq\operatorname{sign}(x)|x|^{\rho},$
where $\operatorname{sign}(x)$ denotes the sign function. As illustrated in Figure 3, the power-uniform discretization with $\rho>1$ induces a more rapid decay in terms of $\eta$ in the high-noise regime (where $\lambda<0$), which corresponds to the early stages of the reverse sampling process.
(a) Steps $N=10$.
(b) Steps $N=20$.
Figure 3: Visualization of the $\eta_{i}=e^{-\lambda_{i}}$ under different time discretization strategies.
Given the discretization in terms of $\lambda$, the discrete time steps $t_{i}$ used in the reverse sampling process (3) are obtained via the inverse mapping $t_{i}=\lambda^{-1}(\lambda_{i}^{(\rho)})$. This mapping relies on the specific definitions of $\alpha(t)$ and $\sigma(t)$. For example, for the cosine schedule defined as $\alpha(t)=\cos\!\left(\frac{\pi}{2}t\right)$ and $\sigma(t)=\sin\!\left(\frac{\pi}{2}t\right)$, this relationship yields the closed-form solution:
$t_{i}=\frac{2}{\pi}\arctan\!\left(e^{-\lambda_{i}^{(\rho)}}\right).$ (82)
Synthetic Source Distribution
Formally, the eigenvalues $\{\mu_{\ell}\}_{\ell=1}^{k}$ of the source covariance matrix are generated using a shifted power-law model. This formulation is designed to simulate the heavy-tailed eigenvalue spectra:
$\mu_{\ell}=C\cdot(\ell+i_{0})^{-p}+\varepsilon,\qquad\ell=1,\ldots,k,$ (83)
where $p>0$ governs the decay rate, $i_{0}>0$ smooths the spectral head. The parameter $C$ is calibrated to satisfy the dominant eigenvalue condition $\mu_{1}=\mu_{\max}$, while the spectral floor $\varepsilon>0$ is tuned to ensure the minimal eigenvalue level $\mu_{k}\geq\mu_{\min}$.
Visualization of Generated Images
In this part, we present additional qualitative comparisons of samples generated by different pretrained models, employing representative ODE samplers across various time discretization strategies. The visual results are provided on the following pages.
Our implementation builds upon the open-source frameworks listed in Table II. We gratefully acknowledge the authors for making their code and pretrained models publicly available.
TABLE II: List of open-source repositories referenced in this work.
Method Source Code URL
DDIM [19] https://github.com/ermongroup/ddim
Score SDE [20] https://github.com/yang-song/score_sde_pytorch
EDM [6] https://github.com/NVlabs/edm
DPM-Solver [12, 13] https://github.com/LuChengTHU/dpm-solver
UniPC [27] https://github.com/wl-zhao/UniPC
DPM-Solver-v3 [28] https://github.com/thu-ml/DPM-Solver-v3
DM-NonUniform [25] https://github.com/scxue/DM-NonUniform
(a) $\rho=1.5$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 4: EDM checkpoint with DPM-Solver++ on CIFAR-10 (NFEs=6).
(a) $\rho=1.5$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 5: EDM checkpoint with DPM-Solver++ on CIFAR-10 (NFEs=12).
(a) $\rho=1.5$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 6: EDM checkpoint with DPM-Solver++ on CIFAR-10 (NFEs=15).
(a) $\rho=2.0$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 7: VP-SDE checkpoint with UniPC on CIFAR-10 (NFEs=20).
(a) $\rho=2.0$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 8: VP-SDE checkpoint with UniPC on CIFAR-10 (NFEs=25).
(a) $\rho=2.0$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 9: VP-SDE checkpoint with UniPC on CIFAR-10 (NFEs=50).
(a) $\rho=1.5$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 10: EDM checkpoint with DPM-Solver++ on FFHQ-64 (NFEs=8).
(a) $\rho=1.5$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 11: EDM checkpoint with DPM-Solver++ on FFHQ-64 (NFEs=10).
(a) $\rho=1.5$.
(b) Uniform-$\lambda$.
(c) Xue’24.
Figure 12: EDM checkpoint with DPM-Solver++ on FFHQ-64 (NFEs=12).

### v2

@@
 An elementary approach to characterizing the impact of noise scheduling and time discretization in generative diffusion models is developed.
-Considering a simplified model where the source distribution is multivariate Gaussian with a given covariance matrix, the explicit closed-form evolution trajectory of the distributions across reverse sampling steps is derived, and consequently, the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output is obtained.
+We first utilize the Cramér-Rao bound to identify the Gaussian setting as a fundamental performance limit, necessitating its study as a reference.
+Building on this insight, we consider a simplified model in which the source distribution is a multivariate Gaussian with a given covariance matrix, together with the deterministic reverse sampling process.
+The explicit closed-form evolution trajectory of the distributions across reverse sampling steps is derived, and consequently, the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output is obtained.
 The effect of the number of time discretization steps on the convergence of this KL divergence is studied via the Euler-Maclaurin expansion.
@@
 However, the majority of extant works rely on computationally expensive data-driven retraining or evaluation processes for searching the optimal steps, typically incurring a substantial overhead.
-Aiming at understanding from first principles, the interplay between noise scheduling and sampling efficiency, in this work, we adopt an elementary approach based on an idealized case where the source distribution is multivariate Gaussian with a given covariance matrix.
-Such a simple model induces a linear form of the optimal posterior estimator, and consequently an explicit closed-form characterization of the evolution trajectory of the distributions across reverse sampling steps.
+Aiming at understanding from first principles, the interplay between noise scheduling and sampling efficiency, in this work, we first utilize the Girsanov theorem and the Cramér-Rao bound to reveal that the KL divergence between the continuous and discretized reverse-time SDEs in the Gaussian setting serves as a strict theoretical lower bound for that of general distributions.
+This finding validates the Gaussian setting as a necessary proxy for parameter optimization in DMs, providing a strong motivation for leveraging the analytical properties of Gaussians to analyze diffusion dynamics.
+Consequently, in the remainder of this paper, we adopt an elementary approach focusing on a multivariate Gaussian source with a given covariance matrix.
+Such a simple model induces a linear form of the optimal posterior estimator, and consequently an explicit closed-form characterization of the evolution trajectory of the distributions across the deterministic reverse sampling steps.
 Using this closed-form formulation, we can conveniently investigate the Kullback-Leibler (KL) divergence between the source distribution and the reverse sampling output.
@@
 We test this approach over CIFAR-10 [12] and FFHQ-64 [10] datasets with different model configurations as outlined in [9, 25], and the experimental results confirm that our strategy consistently demonstrates superior performance in comparison to baseline and state-of-the-art search-based strategies, particularly when the budget on NFEs is very tight.
-Notation: For brevity, we abbreviate a function $f(t)$ as $f$ and denote its discrete-time counterpart $f(t_{j})$ by $f_{t_{j}}$.
+Notation: When there is no ambiguity, we omit the explicit time argument $t$ for time-dependent functions (e.g., writing $f$ for $f(t)$) and denote its discrete-time counterpart $f(t_{j})$ by $f_{t_{j}}$.
 Derivatives with respect to $t$ are written as $\dot{f}=\mathrm{d}f/\mathrm{d}t$ and $\ddot{f}=\mathrm{d}^{2}f/\mathrm{d}t^{2}$.
-Throughout $\log(\cdot)$ represents the natural logarithm.
+Throughout this paper, $\log(\cdot)$ represents the natural logarithm.
 II Preliminaries
-In this section, we first review the mechanism of DMs, encompassing the forward diffusion process with its noise schedule design, and the deterministic reverse sampling process employed for generation.
-Subsequently, we outline the multivariate Gaussian setting that enables a tractable theoretical analysis.
-II-A Diffusion Models
+In this section, we first review the mechanism of DMs, encompassing the forward diffusion process with its noise schedule design, and the reverse sampling process employed for generation.
 The forward diffusion process of DMs is characterized by a linear Gaussian perturbation [8, 11]:
 $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (1)
-where $t\in[0,1]$ denotes the diffusion time, $\mathbf{x}_{0}\sim q(\mathbf{x})$ represents the source data sample, and $\alpha_{t},\sigma_{t}$ parameterize the noise schedule that directly controls the noise-adding process.
-Commonly adopted noise schedules include two types: variance-preserving (VP) and variance-exploding (VE) [20].
-The VP setting usually satisfies the constraint $\alpha^{2}(t)+\sigma^{2}(t)=1$, where $\alpha(t)$ decays from 1 to 0.
-In contrast, the VE setting maintains $\alpha(t)\equiv 1$, while $\sigma(t)$ increases from 0 to a predefined maximum $\sigma_{\max}\gg 1$.
-In either case, the forward transition kernel is given by
-$q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}\!\left(\alpha_{t}\mathbf{x}_{0},\,\sigma^{2}_{t}\mathbf{I}\right),$ (2)
-which progressively diffuses the source distribution $q(\mathbf{x}_{0})$ towards a Gaussian noise as $t\to 1$.
-The deterministic reverse sampling process, which reverses the forward diffusion process (1) to generate data samples from $q(\mathbf{x})$, is governed by the following update formula [19, 12], for each step $j\in\left[1,N\right]$:
-$\hat{\mathbf{x}}_{t_{j-1}}=\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\hat{\mathbf{x}}_{t_{j}}+\left(\sigma_{t_{j-1}}-\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\sigma_{t_{j}}\right)\bm{\epsilon}_{\theta}(\hat{\mathbf{x}}_{t_{j}},t_{j}),$ (3)
-where $\{t_{i}\}_{i=0}^{N}$ denotes the time discretization sequence decreasing from $t_{N}=1$ to $t_{0}=0$, and $\bm{\epsilon}_{\theta}$ is a learned noise estimator.
-The reverse sampling process is initialized with $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ for the VP setting and $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{\max}^{2}\mathbf{I})$ for the VE setting.
-To obtain the noise estimator $\bm{\epsilon}_{\theta}$, the training objective is to minimize the following mean squared error [5]:
+where $t\in[0,1]$ denotes the diffusion time, $\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0})$ represents the source data sample, and $\alpha_{t},\sigma_{t}$ parameterize the noise schedule that directly controls the noise-adding process.
+Commonly adopted noise schedules in DMs are categorized into two types: variance-preserving (VP) and variance-exploding (VE) [25].
+The VP schedule is constrained by $\alpha_{t}^{2}+\sigma_{t}^{2}=1$, with boundary conditions $\lim_{t\to 0}\alpha_{t}=1$ and $\lim_{t\to 1}\alpha_{t}=0$.
+In contrast, the VE schedule maintains $\alpha_{t}\equiv 1$ while satisfying $\lim_{t\to 0}\sigma_{t}=0$ and $\lim_{t\to 1}\sigma_{t}=\sigma_{\max}\gg 1$.
+Despite these differences, the forward transition kernel for both schedules unifies to:
+$q_{0t}(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x}_{0},\sigma_{t}^{2}\mathbf{I}),$ (2)
+which progressively diffuses the source distribution $q_{0}(\mathbf{x}_{0})$ toward a tractable Gaussian prior as $t\to 1$.
+Moreover, the trajectory of the marginal distributions governed by (2) can be equivalently described by a SDE [25, 11]:
+$\mathrm{d}\mathbf{x}_{t}=f(t)\mathbf{x}_{t}\mathrm{d}t+g(t)\mathrm{d}\mathbf{w}_{t},\quad\mathbf{x}_{0}\sim q_{0}(\mathbf{x}_{0}),$ (3)
+where $\mathbf{w}_{t}$ denotes the standard Wiener process. The drift coefficient $f(t)$ and diffusion coefficient $g(t)$ are derived as [11]:
+$f(t)=\frac{\dot{\alpha}}{\alpha},\quad g(t)=\sqrt{2\sigma\left(\dot{\sigma}-\frac{\dot{\alpha}}{\alpha}\sigma\right)}.$ (4)
+According to the classical theory of diffusion processes [1, 25], the corresponding reverse-time SDE flowing from $t=1$ to 0 is given by:
+$\mathrm{d}\mathbf{x}_{t}=\left(f(t)\mathbf{x}_{t}-g^{2}(t)\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})\right)\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}_{t},$ (5)
+where $\bar{\mathbf{w}}_{t}$ is a standard Wiener process in the reverse time.
+Theoretically, when initialized at $\mathbf{x}_{1}\sim q_{1}(\mathbf{x}_{1})$, this process reverses the diffusion trajectory and enables sampling from the source distribution $q_{0}(\mathbf{x}_{0})$.
+To render the reverse-time SDE tractable, DMs employ a neural network $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ to approximate the scaled score function, i.e., $\sigma_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$, which is optimized via the following mean squared error (MSE) objective:
 $\mathcal{L}_{\theta}=\mathbb{E}_{t\in\left[0,1\right],\mathbf{x}_{0}\sim q(\mathbf{x}),\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\mathbf{I}\right)}\!\left[\bigl\|\bm{\epsilon}-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\bigr\|^{2}\right],$ (6)
@@
 Inverting the linear forward relation in (1), then gives the identity:
-$\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=\mathbb{E}\left[\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}}\bigg|\mathbf{x}_{t}\right]=\frac{\mathbf{x}_{t}-\alpha_{t}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\sigma_{t}}.$ (5)
+$\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=\mathbb{E}\left[\frac{\mathbf{x}_{t}-\alpha_{t}\mathbf{x}_{0}}{\sigma_{t}}\bigg|\mathbf{x}_{t}\right]=\frac{\mathbf{x}_{t}-\alpha_{t}\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]}{\sigma_{t}}=-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t}),$ (7)
+where the last equality follows from Tweedie’s formula [4].
 Thus, learning to predict the noise $\bm{\epsilon}$ is mathematically equivalent to estimating the posterior mean of the source data $\mathbf{x}_{0}$.
-II-B Gaussian Setup
-To facilitate theoretical analysis, we consider the idealized case where the source distribution is multivariate Gaussian with a given covariance matrix, i.e., $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\bm{\Sigma}_{\mathbf{x}})$.
+To derive an implementable numerical scheme, we discretize the reverse-time SDE (5).
+Consider a time discretization sequence $\{t_{i}\}_{i=0}^{N}$ decreasing from $t_{N}=1$ to $t_{0}=0$.
+For the integration step from $t_{j}$ to $t_{j-1}$ (i.e., the interval $t\in(t_{j-1},t_{j}]$), we approximate the score component by freezing the neural estimator at the starting time $t_{j}$.
+This yields the following piecewise SDE approximation [2, 15]:
+$\mathrm{d}\mathbf{x}_{t}=\left(f(t)\mathbf{x}_{t}-h(t)\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right)\mathrm{d}t+g(t)\mathrm{d}\bar{\mathbf{w}}_{t},\quad\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}),$ (8)
+where the coefficient $h(t)$ is derived as:
+$h(t)=\frac{g^{2}(t)}{-\sigma(t)}=2\left(\frac{\dot{\alpha}}{\alpha}\sigma-\dot{\sigma}\right).$ (9)
+In practice, the initialization parameter $\sigma_{N}$ is set by the noise schedule, with $\sigma_{N}=1$ for the VP setting and $\sigma_{N}=\sigma_{\max}$ for the VE setting.
+For subsequent theoretical analysis, we denote the path measures induced by the exact reverse-time SDE (5) and the approximate reverse-time SDE (8) as $\mathbb{Q}$ and $\mathbb{P}$, respectively.
+While the stochastic sampling above recovers the data distribution, deterministic sampling is often preferred in practice for its stability and effectiveness.
+By converting the SDE into its corresponding probability flow ODE (PF-ODE) [25], we obtain the deterministic reverse sampling process [24, 16].
+This process inverts the forward perturbation to generate samples from $q_{0}(\mathbf{x}_{0})$ via the following iterative update rule for $j=N,N-1,\dots,1$:
+$\hat{\mathbf{x}}_{t_{j-1}}=\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\hat{\mathbf{x}}_{t_{j}}+\left(\sigma_{t_{j-1}}-\frac{\alpha_{t_{j-1}}}{\alpha_{t_{j}}}\sigma_{t_{j}}\right)\bm{\epsilon}_{\theta}(\hat{\mathbf{x}}_{t_{j}},t_{j}),$ (10)
+where the process is initialized with $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I})$.
+III Gaussian Reference and Lower Bound
+This section provides the theoretical motivation for our focus on the Gaussian setting.
+Specifically, by utilizing stochastic analysis on the reverse-time SDEs, we demonstrate that the Gaussian case serves as a fundamental lower bound for general distributions.
+We first introduce a Gaussian reference process.
+Let $q_{0,G}(\mathbf{x}_{0})$ be a Gaussian distribution constructed to share the identical covariance structure with the underlying data distribution $q_{0}(\mathbf{x})$.
+Specifically, denoting the data covariance as $\mathrm{Cov}_{q_{0}}[\mathbf{x}_{0}]=\bm{\Sigma}_{\mathbf{x}}$, we consider the reference distribution $q_{0,G}$ with covariance $\bm{\Sigma}_{\mathbf{x}}$.
+Due to the linearity of the forward diffusion governed by (1), this covariance matching is preserved throughout the forward process, i.e.
+$\mathrm{Cov}_{q_{t}}[\mathbf{x}_{t}]=\mathrm{Cov}_{q_{t,G}}[\mathbf{x}_{t}]=\bm{\Sigma}_{t}=\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}$.
+Moreover, the Hessian of the log-density for this Gaussian reference is given directly by:
+$\nabla_{\mathbf{x}_{t}}^{2}\log q_{t,G}(\mathbf{x}_{t})=-\bm{\Sigma}_{t}^{-1}.$ (11)
+Analogous to the general case, we define $\mathbb{Q}_{G}$ and $\mathbb{P}_{G}$ as the path measures induced by the exact reverse-time SDE and the approximate reverse-time SDE, respectively, operating under the Gaussian source assumption $q_{0,G}$.
+Theorem 1.
+Under the optimal estimation, the KL divergence of the Gaussian reference process serves as a lower bound for the true process:
+$D_{KL}(\mathbb{Q}\|\mathbb{P})\geq D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})\geq D_{KL}(q_{0,G}\|p_{0,G}).$ (12)
+Proof.
+We first aim to prove the inequality $D_{KL}(\mathbb{Q}\|\mathbb{P})\geq D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})$.
+Let $\Delta$ denote the gap between the KL divergence of the general process and that of the Gaussian reference process.
+Under the assumption that the neural network approximates the optimal score, i.e., $\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=-\sigma_{t}\nabla_{\mathbf{x}_{t}}\log q_{t}(\mathbf{x}_{t})$ given in (7), its Jacobian is given by $\nabla_{\mathbf{x}}\bm{\epsilon}^{*}_{\theta}(\mathbf{x}_{t},t)=-\sigma_{t}\nabla_{\mathbf{x}_{t}}^{2}\log q_{t}(\mathbf{x}_{t})$.
+Specifically for the Gaussian reference process, this simplifies to a constant Jacobian $\nabla_{\mathbf{x}}\bm{\epsilon}^{*}_{\theta,G}(\mathbf{x}_{t},t)=\sigma_{t}\bm{\Sigma}_{t}^{-1}$ due to (11).
+Consequently, using Lemma 3 in Appendix E, the difference decomposes as:
+$\Delta$ $\triangleq D_{KL}(\mathbb{Q}\|\mathbb{P})-D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})$
+$=\Delta_{init}+\sum_{j=1}^{N}\frac{1}{4}h_{t_{j}}^{2}\delta_{j}^{2}\left(\mathbb{E}_{q_{t_{j}}}\left[\|-\sigma_{t_{j}}\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})\|_{F}^{2}\right]-\|\sigma_{t_{j}}\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)+\mathcal{R}(\delta),$ (13)
+where the initial term $\Delta_{init}$ is defined as:
+$\Delta_{init}=D_{KL}(q_{1}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))-D_{KL}(q_{1,G}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I})).$ (14)
+For the initial term $\Delta_{\text{init}}$, we utilize the decomposition of KL divergence: $D_{KL}(q\|p)=-H(q)+H(q,p)$, where $H(q,p)=-\mathbb{E}_{q}[\log p]$ is the cross-entropy.
+Since the prior $p_{\text{prior}}=\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I})$ is Gaussian, the cross-entropy term depends solely on the first two moments of $q$.
+By construction, the Gaussian reference $q_{1,G}$ shares the identical mean and covariance with the true distribution $q_{1}$, we have
+$\mathbb{E}_{q_{1}}[\log p_{\text{prior}}]=\mathbb{E}_{q_{1,G}}[\log p_{\text{prior}}].$ (15)
+The difference $\Delta_{\text{init}}$ thus simplifies to the difference in differential entropies
+$\Delta_{\text{init}}=\left(-H(q_{1})\right)-\left(-H(q_{1,G})\right)=H(q_{1,G})-H(q_{1}).$ (16)
+According to the Maximum Entropy Theorem [3], among all distributions with a fixed covariance matrix, the Gaussian distribution maximizes the differential entropy.
+Therefore, $H(q_{1,G})\geq H(q_{1})$, implying $\Delta_{\text{init}}\geq 0$.
+To determine the sign of $\Delta$, we analyze the Jacobian terms. The difference can be expanded as:
+$\mathbb{E}_{q_{t_{j}}}\left[\|-\sigma_{t_{j}}\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})\|_{F}^{2}\right]-\|\sigma_{t_{j}}\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}$
+$=\sigma_{t_{j}}^{2}\left(\mathbb{E}_{q_{t_{j}}}\left[\|\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})\|_{F}^{2}\right]-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$ (17)
+$\geq\sigma_{t_{j}}^{2}\left(\left\|\mathbb{E}_{q_{t_{j}}}[\nabla_{\mathbf{x}}^{2}\log q_{t_{j}}(\mathbf{x})]\right\|_{F}^{2}-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$ (18)
+$=\sigma_{t_{j}}^{2}\left(\|-\mathcal{I}(q_{t_{j}})\|_{F}^{2}-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$ (19)
+$=\sigma_{t_{j}}^{2}\left(\|\mathcal{I}(q_{t_{j}})\|_{F}^{2}-\|\bm{\Sigma}_{t_{j}}^{-1}\|_{F}^{2}\right)$
+$\geq 0,$ (20)
+where $\mathcal{I}(q_{t})\triangleq\mathbb{E}_{q_{t}}[-\nabla_{\mathbf{x}}^{2}\log q_{t}(\mathbf{x})]$ denotes the Fisher Information Matrix.
+The inequality in (18) follows from Jensen’s inequality applied to the convex function $\|\cdot\|_{F}^{2}$.
+Equality (19) utilizes the identity that the expected Hessian of the log-likelihood is the negative of the Fisher Information Matrix.
+Finally, the inequality in (20) is guaranteed by the multivariate Cramér-Rao Bound, which states that $\mathcal{I}(q_{t})\succeq\bm{\Sigma}_{t}^{-1}$ [3], and all the equalities hold if and only if $q_{t}(\mathbf{x})$ is also a Gaussian distribution.
+Substituting this result and $\Delta_{init}\geq 0$ into (III), we obtain $\Delta\geq 0$ for a sufficiently small step size.
+Furthermore, the second inequality $D_{KL}(\mathbb{Q}_{G}\|\mathbb{P}_{G})\geq D_{KL}(q_{0,G}\|p_{0,G})$ follows directly from the Data Processing Inequality (DPI), as the marginal distribution at $t=0$ is a projection of the path measure.
+∎
+Theorem 1 reveals that the Gaussian setting characterizes a theoretical lower bound of the KL divergence for general distributions.
+This motivates our subsequent analysis of the Gaussian source, serving as a tractable proxy for general distributions, to derive optimal parameter designs for DMs.
+IV KL-Based Analysis on Gaussian Setting and Applications
+In this section, we present the results derived from our theoretical analysis under the Gaussian setting and the widely used deterministic reverse sampling process introduced in (10), and demonstrate their applications.
+For clarity of exposition, we focus on the VP setting; analogous results for the VE setting are briefly described in discussions and remarks.
+As the setup given in Section III, we suppose $\mathbf{x}_{0}\sim q_{0,G}(\mathbf{x}_{0})=\mathcal{N}(\mathbf{0},\bm{\Sigma}_{\mathbf{x}})$, so the marginal distribution $q_{t}(\mathbf{x}_{t})$ remains Gaussian:
+$q_{t,G}(\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{0},\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right).$ (21)
+Moreover, in the Gaussian setting, the optimal posterior estimator admits a closed-form solution [20]:
+$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\alpha_{t}\bm{\Sigma}_{\mathbf{x}}\left(\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right)^{-1}\mathbf{x}_{t}.$ (22)
 Without loss of generality, we assume that the covariance matrix $\bm{\Sigma}_{\mathbf{x}}$ is positive definite.
 Let the eigendecomposition of the covariance matrix be $\bm{\Sigma}_{\mathbf{x}}=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{-1}$, where $\mathbf{U}$ is an orthogonal matrix and $\mathbf{\Lambda}=\text{diag}(\mu_{1},\ldots,\mu_{k})$ denotes the diagonal matrix of eigenvalues.
-Furthermore, due to the linearity of the forward diffusion process (1), the marginal distribution $q(\mathbf{x}_{t})$ remains Gaussian:
-$q(\mathbf{x}_{t})=\mathcal{N}\left(\mathbf{0},\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right).$ (6)
-Moreover, the optimal posterior estimator admits a closed-form solution [15]:
-$\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]=\alpha_{t}\bm{\Sigma}_{\mathbf{x}}\left(\alpha_{t}^{2}\bm{\Sigma}_{\mathbf{x}}+\sigma_{t}^{2}\mathbf{I}\right)^{-1}\mathbf{x}_{t}.$ (7)
-III KL-Based Analysis and Applications
-In this section, we present the main results derived from our theoretical analysis under the Gaussian setting, and demonstrate their applications.
-For clarity of exposition, we focus on the VP setting; analogous results for the VE setting are briefly described in discussions and remarks.
-III-A KL Divergence Analysis
+IV-A KL Divergence Analysis
 We begin by characterizing the evolution trajectory of the distributions across reverse sampling steps, as given by the following lemma.
 Lemma 1.
-With the initialization $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, time discretization sequence $\{t_{i}\}_{i=0}^{N}$, and the optimal estimator given by (5) and (7), the generated sample $\hat{\mathbf{x}}_{t_{0}}$ produced by the reverse sampling process (3) is distributed as
-$p\left(\hat{\mathbf{x}}_{t_{0}}\right)=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),\quad\mathbf{M}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (8)
+With the initialization $\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, time discretization sequence $\{t_{i}\}_{i=0}^{N}$, and the optimal estimator given by (7) and (22), the generated sample $\hat{\mathbf{x}}_{t_{0}}$ produced by deterministic reverse sampling (10) is distributed as
+$\hat{p}_{0,G}\left(\hat{\mathbf{x}}_{t_{0}}\right)=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),\quad\mathbf{M}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (23)
 where the $\ell$-th eigenvalue is given by
@@
 See Appendix A. ∎
-By (6), we can rewrite $q(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$, where $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$ with
+By (21), we can rewrite $q_{0,G}(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$, where $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$ with
 $n_{\ell}\triangleq\alpha_{t_{0}}^{2}\mu_{\ell}+\sigma_{t_{0}}^{2}.$ (25)
 According to Lemma 1, the KL divergence is expressible in closed form:
-$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)=\frac{1}{2}\sum_{\ell=1}^{k}\left(\frac{m_{\ell}}{n_{\ell}}-\log\frac{m_{\ell}}{n_{\ell}}-1\right),$ (11)
+$D_{\mathrm{KL}}\!\left(\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})\,\|\,q_{0,G}(\mathbf{x}_{t_{0}})\right)=\frac{1}{2}\sum_{\ell=1}^{k}\left(\frac{m_{\ell}}{n_{\ell}}-\log\frac{m_{\ell}}{n_{\ell}}-1\right),$ (26)
 by the standard KL divergence expression between two Gaussian distributions.
@@
 This implies that $r_{\ell}=\mathcal{O}(1/N)$, which justifies the use of the Taylor expansion for large $N$. Substituting (32) into (26), the KL divergence can be expanded as follows:
-$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)$ $=\frac{1}{2}\sum_{\ell=1}^{k}\bigl(e^{2r_{\ell}}-2r_{\ell}-1\bigr)$
+$D_{\mathrm{KL}}\!\left(\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})\,\|\,q_{0,G}(\mathbf{x}_{t_{0}})\right)$ $=\frac{1}{2}\sum_{\ell=1}^{k}\bigl(e^{2r_{\ell}}-2r_{\ell}-1\bigr)$
 $=\sum_{\ell=1}^{k}r_{\ell}^{2}+\mathcal{O}\!\left(\sum_{\ell=1}^{k}|r_{\ell}|^{3}\right)$
@@
 From (33), we observe that the KL divergence scales as $\mathcal{O}(1/N^{2})$.
-As $N\to\infty$, the eigenvalue mismatch vanishes (i.e., $m_{\ell}/n_{\ell}\to 1$), leading to $D_{\mathrm{KL}}(p(\hat{\mathbf{x}}_{t_{0}})\|q(\mathbf{x}_{t_{0}}))\to 0$.
+As $N\to\infty$, the eigenvalue mismatch vanishes (i.e., $m_{\ell}/n_{\ell}\to 1$), leading to $D_{\mathrm{KL}}(p_{0,G}(\hat{\mathbf{x}}_{t_{0}})\|q_{0,G}(\mathbf{x}_{t_{0}}))\to 0$.
 This verifies that the reverse sampling process in (10) is asymptotically consistent, recovering the exact source distribution in the continuous-time limit, in accordance with the classical theory well known for DMs (see, e.g., [1]).
-III-B Derivation of Tangent Law
+IV-B Derivation of Tangent Law
 The preceding analysis of the KL divergence reveals a direct connection between the noise schedule and sampling accuracy, culminating in the formulation of the optimization problem (34).
@@
 $\min_{\eta}\mathcal{J}_{\ell}[\eta]\triangleq\int_{0}^{1}\frac{\dot{\eta}^{2}}{\bigl(\mu_{\ell}+\eta^{2}\bigr)^{2}}\,\mathrm{d}t,$
-$\text{subject to }\eta(0)=0\text{ and }\lim_{t\to 1}\eta(t)=\infty.$ (21)
+$\text{subject to }\lim_{t\to 0}\eta(t)=0\text{ and }\lim_{t\to 1}\eta(t)=\infty.$ (36)
 The unique solution $\eta(\cdot)$ that minimizes the functional $\mathcal{J}_{\ell}$ is given by the tangent law schedule $\eta_{\ell}(t)$ in Definition 1.
@@
 implying that the optimal $Q_{\ell}(t)$ is linear in $t$: $Q_{\ell}(t)=ct+d$.
-Given the boundary conditions $\eta(0)=0$ and $\lim_{t\to 1}\eta(t)=\infty$, the transformed boundary values are $Q_{\ell}(0)=0$ and $Q_{\ell}(1)=\frac{\pi}{2\sqrt{\mu_{\ell}}}$.
+Given the boundary conditions in (36), the transformed boundary values are $Q_{\ell}(0)=0$ and $Q_{\ell}(1)=\frac{\pi}{2\sqrt{\mu_{\ell}}}$.
 Therefore, the optimal trajectory is $Q_{\ell}(t)=\frac{\pi t}{2\sqrt{\mu_{\ell}}}$.
@@
 Additional numerical results demonstrating that the parameterized tangent law schedule outperforms other noise schedules for the Gaussian setting are presented in Appendix D-B.
-III-C Time Discretization Strategy Evaluation
+IV-C Time Discretization Strategy Evaluation
 In this subsection, we demonstrate another more direct application of the KL divergence-based analysis.
@@
 To apply (26), we model the source distribution using a synthetic power-law covariance spectrum, which mimics the heavy-tailed eigenvalue decay characteristic of real image data [26].
-Figure 1 presents the quantitative evaluation of the closed-form KL divergence (11) across varying step budgets $N$ (details in Appendix E).
+Figure 1 presents the quantitative evaluation of the closed-form KL divergence (26) across varying step budgets $N$ (details in Appendix F).
 This comparison covers both VP and VE settings as considered in [25] with the same time discretization strategy given in (45).
@@
 These preliminary observations motivate further experimentation on real-world image generation tasks, as presented in the next section.
-IV Experiments
+V Experiments
 Focusing on the practical image generation tasks using pretrained DMs, we evaluate the performance of our selected time discretization strategies in conjunction with off-the-shelf samplers.
@@
 Conversely, our approach maintains consistent superiority across all budgets without incurring any costly overhead, validating its effectiveness across varying model architectures and datasets.
-V Conclusion
+VI Conclusion
 This work has presented a first-principles analysis of diffusion sampling, establishing an explicit connection between the noise schedule and sampling accuracy.
@@
 By (21), in the practical reverse sampling, the reverse process is initialized from the terminal prior:
-$\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{t_{N}}^{2}\mathbf{I}),\quad\text{where}\quad\sigma_{t_{N}}=\begin{cases}1&\text{VP setting},\\ \sigma_{\max}&\text{VE setting}.\end{cases}$ (37)
+$\hat{\mathbf{x}}_{t_{N}}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}),\quad\text{where}\quad\sigma_{N}=\begin{cases}1&\text{VP setting},\\ \sigma_{\max}&\text{VE setting}.\end{cases}$ (52)
 Since the update (51) is affine transformation on a Gaussian distribution, $\hat{\mathbf{x}}_{t_{0}}$ remains Gaussian:
-$p(\hat{\mathbf{x}}_{t_{0}})=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),$ (38)
+$\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})=\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right),$ (53)
 where the diagonal matrix $\mathbf{M}$ is
-$\mathbf{M}$ $\triangleq\sigma_{t_{N}}^{2}\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)^{\!\top}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (39)
+$\mathbf{M}$ $\triangleq\sigma_{N}^{2}\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)\left(\prod_{j=1}^{N}\mathbf{D}_{j}\right)^{\!\top}=\mathrm{diag}(m_{1},\ldots,m_{k}),$ (54)
 with its diagonal entries
-$m_{\ell}=\sigma_{t_{N}}^{2}\prod_{j=1}^{N}d_{j,\ell}^{2},\qquad\ell=1,\ldots,k,$ (40)
+$m_{\ell}=\sigma_{N}^{2}\prod_{j=1}^{N}d_{j,\ell}^{2},\qquad\ell=1,\ldots,k,$ (55)
 where $d_{j,\ell}$ is given in (50). This yields the result stated in Lemma 1.
@@
 Finally, combining the distribution $q(\mathbf{x}_{t_{0}})=\mathcal{N}(\mathbf{0},\mathbf{U}\mathbf{N}\mathbf{U}^{-1})$ with $\mathbf{N}=\mathrm{diag}(n_{1},\ldots,n_{k})$, we can immediately obtain
-$D_{\mathrm{KL}}\!\left(p(\hat{\mathbf{x}}_{t_{0}})\,\|\,q(\mathbf{x}_{t_{0}})\right)$ $=D_{\mathrm{KL}}\!\left(\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right)\,\|\,\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{N}\mathbf{U}^{-1}\right)\right)$
+$D_{\mathrm{KL}}\!\left(\hat{p}_{0,G}(\hat{\mathbf{x}}_{t_{0}})\,\|\,q_{0,G}(\mathbf{x}_{t_{0}})\right)$ $=D_{\mathrm{KL}}\!\left(\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{M}\mathbf{U}^{-1}\right)\,\|\,\mathcal{N}\!\left(\mathbf{0},\,\mathbf{U}\mathbf{N}\mathbf{U}^{-1}\right)\right)$
 $=\frac{1}{2}\left(\mathrm{tr}(\mathbf{N}^{-1}\mathbf{M})-\log\det(\mathbf{N}^{-1}\mathbf{M})-k\right)$
@@
 To interpret this result, we first recall that $n_{\ell}\triangleq\alpha^{2}(0)\mu_{\ell}+\sigma^{2}(0)$ (25).
-We now consider the following two boundary conditions, as we mentioned before in Section II-A.
-  • Variance Preserving (VP): The boundary conditions are specified as $\alpha(0)=1,\sigma(0)=0$ and $\alpha(1)=0,\sigma(1)=1$. Substituting these into (61) yields
+We now consider the following two boundary conditions, as we mentioned before in Section II.
+  • Variance Preserving (VP): The boundary conditions are specified as $\lim_{t\to 0}\alpha(t)=1$, $\lim_{t\to 0}\sigma(t)=0$, and typically $\lim_{t\to 1}\alpha(t)=0$, $\lim_{t\to 1}\sigma(t)=1$.
+Substituting these into (76) yields
     $I_{\ell}\big|_{\text{VP}}=\frac{1}{2}\log\left(\frac{n_{\ell}}{0\cdot\mu_{\ell}+1}\right)=\frac{1}{2}\log n_{\ell}.$ (77)
-  • Variance Exploding (VE): The boundary conditions are $\alpha(t)\equiv 1$, $\sigma(0)=0$, and $\sigma(1)=\sigma_{\max}$. In this case, the integral (61) becomes
+  • Variance Exploding (VE): The boundary conditions are $\alpha(t)\equiv 1$, $\lim_{t\to 0}\sigma(t)=0$, and $\lim_{t\to 1}\sigma(t)=\sigma_{\max}$. In this case, the integral (76) becomes
     $I_{\ell}\big|_{\text{VE}}=\frac{1}{2}\log\!\left(\frac{n_{\ell}}{\mu_{\ell}+\sigma_{\max}^{2}}\right).$ (78)
@@
 While the objective functional remains identical to that of the VP setting (36), the primary difference lies in the boundary condition at $t=1$.
-Given the noise schedule $\alpha(t)\equiv 1$, $\sigma(0)=0$, and $\sigma(1)=\sigma_{\max}$ for the VE setting, along with the definition $\eta(t)=\sigma(t)/\alpha(t)$, we derive the boundary conditions as follows:
-$\eta(0)=0\quad\text{and}\quad\eta(1)=\sigma_{\max}.$ (71)
+Given the noise schedule $\alpha(t)\equiv 1$, $\lim_{t\to 0}\sigma(t)=0$, and $\lim_{t\to 1}\sigma(t)=\sigma_{\max}$ for the VE setting, along with the definition $\eta(t)=\sigma(t)/\alpha(t)$, we derive the boundary conditions as follows:
+$\lim_{t\to 0}\eta(t)=0\quad\text{and}\quad\lim_{t\to 1}\eta(t)=\sigma_{\max}.$ (86)
 Based on the definition $Q_{\ell}(t)\triangleq\frac{1}{\sqrt{\mu_{\ell}}}\arctan\!\left(\frac{\eta(t)}{\sqrt{\mu_{\ell}}}\right)$ from (37), the transformed boundary conditions become
@@
 We specify the boundary values $\eta_{1}$ for the following two settings:
-  • Variance Preserving (VP): Since $\alpha(1)=0$ and $\sigma(1)=1$, we have $\eta_{1}\to\infty$. Consequently, the angle becomes
+  • Variance Preserving (VP): Since $\lim_{t\to 1}\alpha(t)=0$, $\lim_{t\to 1}\sigma(t)=1$, we have $\eta_{1}\to\infty$. Consequently, the angle becomes
     $\Theta_{\gamma}\big|_{\text{VP}}=\frac{\pi}{2}.$
-  • Variance Exploding (VE): Since $\alpha(t)\equiv 1$ and $\sigma(1)=\sigma_{\max}$, we have $\eta_{1}=\sigma_{\max}$,
+  • Variance Exploding (VE): Since $\alpha(t)\equiv 1$ and $\lim_{t\to 1}\sigma(t)=\sigma_{\max}$, we have $\eta_{1}=\sigma_{\max}$,
     $\Theta_{\gamma}\big|_{\text{VE}}=\arctan\!\left(\frac{\sigma_{\max}}{\sqrt{\gamma}}\right)=\frac{\pi}{2}-\arctan\!\left(\frac{\sqrt{\gamma}}{\sigma_{\max}}\right).$
-Same as we mentioned before in Appendix B-A and Appendix C, under the condition of $\sigma_{\max}^{2}\gg\gamma$, the VP setting solution can be regarded as an accurate approximation of the VE setting.
+Same as we stated before in Appendix B-A and Appendix C, under the condition of $\sigma_{\max}^{2}\gg\gamma$, the VP setting solution can be regarded as an accurate approximation of the VE setting.
 Substituting the parameterized tangent law schedule $\eta_{\gamma}(t)=\sqrt{\gamma}\tan(\Theta_{\gamma}t)$ into the expression of $E_{\ell}^{1}$ given in (85), we proceed as follows:
@@
 $=-\frac{\mu_{\ell}\gamma\Theta_{\gamma}}{2}\int_{0}^{\Theta_{\gamma}}\frac{1}{\bigl(\mu_{\ell}+(\gamma-\mu_{\ell})\sin^{2}x\bigr)^{2}}\,\mathrm{d}x.$ (92)
-The standard integral result from [2, formula 2.563.1 and 2.562.1 in pp.177] is
+The standard integral result from [5, formula 2.563.1 and 2.562.1 in pp.177] is
 $\int\frac{\mathrm{d}x}{(a+b\sin^{2}x)^{2}}$ $=\frac{1}{2a(a+b)}\left[(2a+b)\int\frac{\mathrm{d}x}{a+b\sin^{2}x}+\frac{b\sin x\cos x}{a+b\sin^{2}x}\right]$
@@
 First, we compare our parameterized tangent law schedule (using $\gamma^{\star}$) against representative noise schedules widely employed in DMs (Figure 2 top row).
-The results indicate that our method achieves faster convergence in both low-dimensional and high-dimensional synthetic Gaussian sources (details in Appendix E).
+The results indicate that our method achieves faster convergence in both low-dimensional and high-dimensional synthetic Gaussian sources (details in Appendix F).
 Second, we perform an ablation study on the hyperparameter $\gamma$ (Figure 2 bottom row).
@@
 The results demonstrate the effectiveness of our derived $\gamma^{\star}$ in Theorem 3.
-Appendix E Implementation Details
-In this appendix, we provide implementation details for the experiments discussed in Section III-C and Section IV, along with more qualitative results and visual samples.
+Appendix E Lemma of Girsanov’s Theorem
+Lemma 3.
+Let $\mathbb{P}_{\theta}$ denote the path measure induced by the continuous-time SDE parameterized by the neural network $\bm{\epsilon}_{\theta}$, and let $\tilde{\mathbb{P}}_{\theta}$ denote the measure induced by its discretization scheme:
+$\mathbb{P}_{\theta}:\quad$ $\mathrm{d}\mathbf{x}_{t}=\left(f_{t}\mathbf{x}_{t}-h_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\mathrm{d}t+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t},\quad\mathbf{x}_{1}\sim q_{1}(\mathbf{x}_{1}),$ (96a)
+$\tilde{\mathbb{P}}_{\theta}:\quad$ $\mathrm{d}\mathbf{x}_{t}=\left(f_{t}\mathbf{x}_{t}-h_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right)\mathrm{d}t+g_{t}\mathrm{d}\bar{\mathbf{w}}_{t},\quad\mathbf{x}_{1}\sim\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}).$ (96b)
+The KL divergence between these two processes satisfies:
+$D_{KL}(\mathbb{P}_{\theta}||\tilde{\mathbb{P}}_{\theta})$ $=D_{KL}(q_{1}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\frac{1}{2}\int_{0}^{1}\frac{h_{t}^{2}}{g_{t}^{2}}\,\mathbb{E}_{\mathbf{x}_{t}\sim p_{t}}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]\mathrm{d}t$ (97)
+$=D_{KL}(q_{1}(\mathbf{x}_{1})||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\sum_{j=1}^{N}\frac{1}{4}h_{t_{j}}^{2}\delta_{j}^{2}\,\mathbb{E}_{\mathbf{x}_{t_{j}}\sim p_{t_{j}}}\left[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}\right]+\mathcal{R}(\delta),$ (98)
+where $\delta_{j}=t_{j}-t_{j-1}$ is the step size, $\mathbf{J}_{\bm{\epsilon}}\triangleq\nabla_{\mathbf{x}}\bm{\epsilon}_{\theta}(\mathbf{x},t)$ denotes the Jacobian matrix of $\bm{\epsilon}_{\theta}$, and $\mathcal{R}(\delta)$ represents the cumulative higher-order error terms satisfying $\mathcal{R}(\delta)=\mathcal{O}(\delta^{2})$ as $\delta\triangleq\max_{j}\delta_{j}\to 0$.
+Proof.
+The proof proceeds in two steps.
+First, we apply Girsanov’s theorem to derive the exact integral form of the KL divergence.
+Second, we analyze the local discretization error using Itô’s calculus to obtain the leading-order expansion.
+To facilitate the rigorous derivation of the error bound, we first rely on the following regularity conditions. These assumptions are standard in the theoretical analysis of DMs:
+  • Condition 1 (Coefficients): The SDE coefficients $f_{t}$ and $g_{t}$ are Lipschitz continuous with respect to time $t$ and bounded on $[0,1]$.
+  • Condition 2 (Neural Network): The neural estimator $\bm{\epsilon}_{\theta}(\mathbf{x},t)$ is twice continuously differentiable ($C^{2}$) with respect to $\mathbf{x}$ and $C^{1}$ with respect to $t$.
+Moreover, both the network output and its Jacobian are bounded.
+  • Condition 3 (Bounded Moments): The drift term of the score evolution (defined later in Eq.
+(E)) possesses a bounded second moment, i.e., $\sup_{t\in[0,1]}\mathbb{E}[\|\bm{\phi}(\mathbf{x}_{t},t)\|^{2}]<\infty$.
+Let $u_{t}$ denote the difference between the drift terms of $\mathbb{P}_{\theta}$ and $\hat{\mathbb{P}}_{\theta}$, normalized by the diffusion coefficient $g_{t}$. We define:
+$u_{t}\triangleq\frac{-h_{t}\left(\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right)}{g_{t}}.$ (99)
+To ensure the validity of the measure change, we assume Novikov’s condition holds, which requires that the following expectation is finite:
+$\mathbb{E}\left[\exp\left(\frac{1}{2}\int_{0}^{1}\|u_{t}\|^{2}\mathrm{d}t\right)\right]=\mathbb{E}\left[\exp\left(\frac{1}{2}\int_{0}^{1}\frac{h_{t}^{2}}{g_{t}^{2}}\left\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\right\|^{2}\mathrm{d}t\right)\right]<\infty.$ (100)
+Under conditions 1 and 2, Girsanov’s theorem [19] applies. The KL divergence decomposes into the divergence of the initial distributions and the expected path integral of the squared norm of $u_{t}$:
+$D_{KL}(\mathbb{P}_{\theta}||\tilde{\mathbb{P}}_{\theta})=D_{KL}(q_{1}||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\frac{1}{2}\int_{0}^{1}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t.$ (101)
+Substituting the definition of $u_{t}$ back into Eq. (101) yields the equality given in (97).
+We focus on the expectation term within the interval $t\in(t_{j-1},t_{j}]$. Applying Itô’s Lemma [19] to $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$, its dynamics follow:
+$\mathrm{d}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ $=\left(\frac{\partial\bm{\epsilon}_{\theta}}{\partial t}+\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t},t)\left(f_{t}\mathbf{x}_{t}-h_{t}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\frac{1}{2}g_{t}^{2}\nabla_{\mathbf{x}}^{2}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)\mathrm{d}t+g_{t}\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t},t)\mathrm{d}\bar{\mathbf{w}}_{t}$
+$\triangleq\bm{\phi}(\mathbf{x}_{t},t)\mathrm{d}t+g_{t}\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t},t)\mathrm{d}\bar{\mathbf{w}}_{t},$ (102)
+where $\mathbf{J}_{\bm{\epsilon}}=\nabla_{\mathbf{x}}\bm{\epsilon}_{\theta}\in\mathbb{R}^{d\times d}$ is the Jacobian matrix and $\nabla_{\mathbf{x}}^{2}$ denotes the vector Laplacian operator.
+Integrating from $t$ to $t_{j}$:
+$\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})-\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)=\int_{t}^{t_{j}}\bm{\phi}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau+\int_{t}^{t_{j}}g_{\tau}\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{\tau},\tau)\mathrm{d}\bar{\mathbf{w}}_{\tau}.$ (103)
+Then utilizing the Itô Isometry property [19], we get
+$\mathbb{E}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]$ $=\int_{t}^{t_{j}}g_{\tau}^{2}\mathbb{E}\left[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{\tau},\tau)\|_{F}^{2}\right]\mathrm{d}\tau+\mathbb{E}\left[\left\|\int_{t}^{t_{j}}\bm{\phi}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau\right\|^{2}\right].$ (104)
+Assuming the time discretization is sufficiently fine (i.e., the step size $\delta_{j}\triangleq t_{j}-t_{j-1}$ is small enough), we can apply local expansions based on Conditions 1, 2, and 3:
+  • Using the Lipschitz continuity of the coefficients (Condition 1), we approximate the ratio $\frac{h_{t}^{2}}{g_{t}^{2}}$ by its value at $t_{j}$: $\frac{h_{t}^{2}}{g_{t}^{2}}=\frac{h_{t_{j}}^{2}}{g_{t_{j}}^{2}}+\mathcal{O}(t_{j}-t).$
+  • Let $\Phi(\tau)\triangleq g_{\tau}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{\tau},\tau)\|_{F}^{2}]$.
+Under Conditions 1 and 2, $\Phi(\tau)$ is Lipschitz continuous.
+Expanding around $t_{j}$: $\Phi(\tau)=\Phi(t_{j})+\mathcal{O}(t_{j}-\tau)$.
+  • Since the drift $\bm{\phi}$ has a bounded second moment (Condition 3), applying the Cauchy-Schwarz inequality yields:
+    $\mathbb{E}\left[\left\|\int_{t}^{t_{j}}\bm{\phi}(\mathbf{x}_{\tau},\tau)\mathrm{d}\tau\right\|^{2}\right]\leq(t_{j}-t)\int_{t}^{t_{j}}\mathbb{E}\left[\|\bm{\phi}\|^{2}\right]\mathrm{d}\tau=\mathcal{O}((t_{j}-t)^{2}).$ (105)
+Substituting these estimates back into the local error expectation (104), we obtain
+$\mathbb{E}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]$ $=\int_{t}^{t_{j}}\left(\Phi(t_{j})+\mathcal{O}(t_{j}-\tau)\right)\mathrm{d}\tau+\mathcal{O}((t_{j}-t)^{2})$
+$=\Phi(t_{j})(t_{j}-t)+\int_{t}^{t_{j}}\mathcal{O}(t_{j}-\tau)\mathrm{d}\tau+\mathcal{O}((t_{j}-t)^{2})$
+$=g_{t_{j}}^{2}\mathbb{E}_{\mathbf{x}_{t_{j}}}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}](t_{j}-t)+\mathcal{O}((t_{j}-t)^{2}).$ (106)
+Now, we substitute Eq. (E) back into the integral term given in (101). Focusing on the contribution from the $j$-th interval $(t_{j-1},t_{j}]$, the term involving $u_{t}$ becomes
+$\int_{t_{j-1}}^{t_{j}}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t$ $=\int_{t_{j-1}}^{t_{j}}\frac{h_{t}^{2}}{g_{t}^{2}}\mathbb{E}\left[\|\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\bm{\epsilon}_{\theta}(\mathbf{x}_{t_{j}},t_{j})\|^{2}\right]\mathrm{d}t$
+$=\int_{t_{j-1}}^{t_{j}}\left(\frac{h_{t_{j}}^{2}}{g_{t_{j}}^{2}}+\mathcal{O}(t_{j}-t)\right)\left(g_{t_{j}}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}](t_{j}-t)+\mathcal{O}((t_{j}-t)^{2})\right)\mathrm{d}t$
+$=\int_{t_{j-1}}^{t_{j}}\left(h_{t_{j}}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}](t_{j}-t)+\mathcal{O}((t_{j}-t)^{2})\right)\mathrm{d}t.$ (107)
+So we have the integral of the squared drift difference
+$\int_{0}^{1}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t$ $=\sum_{j=1}^{N}\int_{t_{j-1}}^{t_{j}}\mathbb{E}\left[\|u_{t}\|^{2}\right]\mathrm{d}t$
+$=\sum_{j=1}^{N}\left(h_{t_{j}}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}]\int_{t_{j-1}}^{t_{j}}(t_{j}-t)\mathrm{d}t+\int_{t_{j-1}}^{t_{j}}\mathcal{O}((t_{j}-t)^{2})\mathrm{d}t\right)$
+$=\sum_{j=1}^{N}\left(\frac{1}{2}h_{t_{j}}^{2}\delta_{j}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}(\mathbf{x}_{t_{j}},t_{j})\|_{F}^{2}]+\mathcal{O}(\delta_{j}^{3})\right).$ (108)
+Finally, substituting this back into the Girsanov formula (101), and summing over all intervals $j=1,\dots,N$
+$D_{KL}(\mathbb{P}_{\theta}||\tilde{\mathbb{P}}_{\theta})$ $=D_{KL}(q_{1}||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\frac{1}{2}\sum_{j=1}^{N}\left(\frac{1}{2}h_{t_{j}}^{2}\delta_{j}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}\|_{F}^{2}]+\mathcal{O}(\delta_{j}^{3})\right)$
+$=D_{KL}(q_{1}||\mathcal{N}(\mathbf{0},\sigma_{N}^{2}\mathbf{I}))+\sum_{j=1}^{N}\frac{1}{4}h_{t_{j}}^{2}\delta_{j}^{2}\mathbb{E}[\|\mathbf{J}_{\bm{\epsilon}}\|_{F}^{2}]+\mathcal{R}(\delta),$ (109)
+where $\mathcal{R}(\delta)$ represents the higher-order terms, which satisfies $\mathcal{R}(\delta)\triangleq\mathcal{O}((\max_{j}\delta_{j})^{2})$ as $\delta=\max_{j}\delta_{j}\to 0$.
+This recovers the exact form stated in (98).
+∎
+Appendix F Implementation Details
+In this appendix, we provide implementation details for the experiments discussed in Section IV-C and Section V, along with more qualitative results and visual samples.
 Time Discretization Strategy